### v2

Abstract
Directed acyclic graphs (DAGs) are fundamental structures used across many scientific fields. A key concept in DAGs is the least common ancestor (LCA), which plays a crucial role in understanding hierarchical relationships. Surprisingly little attention has been given to DAGs that admit a unique LCA for every subset of their vertices. Here, we characterize such global lca-DAGs and provide multiple structural and combinatorial characterizations. We show that global lca-DAGs have a close connection to join semi-lattices and establish a connection to forbidden topological minors. In addition, we introduce a constructive approach to generating global lca-DAGs and demonstrate that they can be recognized in polynomial time. We investigate their relationship to clustering systems and other set systems derived from the underlying DAGs.
Keywords: Cluster and descendant systems, join semilattice, poset, topological minor, tree-leaf-child DAG, Galled-tree and level-1 networks
1 Introduction
Directed acyclic graphs (DAGs) are widely used models in many fields of application, ranging from algorithms and data structures [1, 32, 10], machine learning and artificial intelligence [36, 27], to biology and bioinformatics [11, 23, 33], operations research and decision making [37], social and economic sciences [13, 2], and transportation and logistics [2]. Each DAG $G=(V,E)$ naturally gives rise to a partially ordered set $(V,\preceq)$, defined by $u\preceq v$ for vertices $u,v\in V$ whenever there is a directed path from $v$ to $u$ in $G$. In particular, each vertex $v\in V$ is naturally associated with the set $\mathtt{D}(v)\coloneqq\{u\in V(G)\mid u\preceq v\}$ of its descendant vertices, i.e., those reachable from $v$. The set of leaf-descendants of $v$ is then defined by $\operatorname{\mathtt{C}}(v)=\mathtt{D}(v)\cap L(G)$ where $L(G)$ is the set of all $\preceq$-minimal vertices, i.e., the leaves in $G$. A least common ancestor (LCA) of a subset $A\subseteq V(G)$ in a DAG $G$ is a $\preceq$-minimal vertex $v$ in $G$ such that $A\subseteq\mathtt{D}(v)$.
The existence of a unique LCA for subsets $A\subseteq V$ plays a central role in numerous applications. In evolutionary networks $G$, for example, a unique LCA of $A$ represents a well-defined ancestral state from which the taxa in $A$ have diverged, and is therefore crucial for the interpretation of evolutionary relationships [22, 30, 34]. Further applications arise in the identification of the most direct common cause in Bayesian networks [36], in object inheritance within programming languages [3], and in algorithms on lattices [25, 3]; see [35] for an overview.
Much of the existing structural theory is concerned with leaf-based LCA questions, which naturally arise in particular in phylogenetic applications [23]. In this context, rooted trees and networks encode hypothetical evolutionary histories, while empirical data are almost exclusively confined to extant species, i.e., the leaves. It is therefore natural that many structural results are formulated in terms of subsets $A\subseteq L(G)$ and the associated leaf-descendant clusters, LCAs, or related common-ancestor concepts [33, 23, 22, 30, 21, 14, 24]. For instance, [39] studies conditions ensuring unique LCAs for leaf subsets of cardinality at most $k$ and relates them to $k$-weak hierarchies, and in [21] classes of DAGs have been characterized in which every non-empty subset of leaves admit a unique LCA.
In contrast, the present work is concerned with the genuinely global question: which DAGs admit a unique LCA for every non-empty subset of vertices, not only for subsets of leaves? This distinction is essential: leaf-based conditions do not in general determine the corresponding vertex-based LCA structure. Related computational problems have been studied for arbitrary vertex subsets, in particular for pairs of vertices: numerous results address the task of determining an element of $\operatorname{LCA}(A)$ when $A\subseteq V(G)$ consists of $|A|=2$ vertices [7, 16, 31]. Generalizations such as deciding whether two vertices admit a unique LCA and computing the complete set of LCAs for a given pair are discussed in [12, 28, 6], and extensions to subsets of size $|A|\geq 3$ have also been considered [35, 41, 29]. Our main goal here is complementary to these algorithmic questions: we characterize those DAGs for which the LCA is uniquely defined for all non-empty subsets of $V(G)$.
This contribution is organized as follows. After introducing notation and preliminary facts, we derive in Section 3 the order-theoretic characterization via join-semilattices and prove that the global lca-property is equivalent to its pairwise version. Section 4 provides a characterization in terms of topological minors. In Section 5, we give a constructive characterization showing that every global lca-DAG can be built recursively from a single-vertex graph. Section 6 reformulates the global lca-property in terms of the leaf-extended DAG $\operatorname{lxt}(G)$. In Section 7, we characterize global lca-DAGs through closedness of descendant, ancestor, and intermediary set systems. Section 8 shows how a DAG can be reconstructed, up to shortcuts, from a canonical Hasse-diagram representation of an associated set system. We conclude in Section 9 with a brief summary and open questions.
2 Notation and preliminary results
Sets
All sets considered here are assumed to be finite. The set $2^{X}$ denotes the powerset of a set $X$. A set system $\mathfrak{C}$ (on $X$) is a subset $\mathfrak{C}\subseteq 2^{X}$. A set system $\mathfrak{C}$ on $X$ is grounded if $\{x\}\in\mathfrak{C}$ for all $x\in X$ and $\emptyset\notin\mathfrak{C}$. A set system $\mathfrak{C}$ on $X$ is rooted if $X\in\mathfrak{C}$. A rooted and grounded set system is a clustering system.
Most of the set system considered here (e.g. those derived from DAGs) will not contain the empty set. Therefore, closedness is naturally defined in terms of non-empty intersections:
Definition 2.1.
A set system $\mathfrak{C}$ is closed if and only if $A,B\in\mathfrak{C}$ implies $A\cap B\in\mathfrak{C}$ or $A\cap B=\emptyset$.
In the theory of convex structures [40], unlike here, set systems are assumed to contain the empty set. This leads to a few subtle, but rather trivial, differences in the definition of closure and closedness that concern the role of the empty set. For instance, including the empty set in $\mathfrak{C}$ relieves the need to treat $A\cap B=\emptyset$ as a separate case in Definition 2.1. We defer this point to at the end of Section 6, see also [9].
[…]
DAGs
A directed graph $G=(V,E)$ is an ordered pair consisting of a non-empty set $V(G)\coloneqq V$ of vertices and a set $E(G)\coloneqq E\subseteq\left(V\times V\right)\setminus\{(v,v)\mid v\in V\}$ of edges. For directed graphs $G=(V_{G},E_{G})$ and $H=(V_{H},E_{H})$, an isomorphism between $G$ and $H$ is a bijective map $\varphi\colon V_{G}\to V_{H}$ such that $(u,v)\in E_{G}$ if and only if $(\varphi(u),\varphi(v))\in E_{H}$. If such a map exists, then $G$ and $H$ are isomorphic, in symbols $G\simeq H$.
A subgraph of $G=(V,E)$ is a directed graph $H=(V^{\prime},E^{\prime})$ such that $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. We occasionally denote this by $H\subseteq G$. If $E^{\prime}$ comprises all edges $(u,v)$ in $E$ with $u,v\in V^{\prime}$, then $H$ is the subgraph induced by $V^{\prime}$. In particular, GvG-v denotes the subgraph of $G$ induced by $V\setminus\{v\}$, provided that $v\in V$ and $|V|>1$. A directed $v_{1}v_{n}$-path $P=(V,E)$ has vertex set $V=\{v_{1},v_{2},\ldots,v_{n}\}$ and the edges in $E$ are precisely of the form $(v_{i},v_{i+1})$, $i=1,2,\ldots,n-1$. If there is an edge $(v_{n},v_{1})$ and a directed $v_{1}v_{n}$-path in $G$, then $G$ contains a cycle. A directed graph without cycles is a directed acyclic graph (DAG).
For a directed graph $G=(V,E)$, we define $\operatorname{indeg}_{G}(v)\coloneqq\left|\left\{u\in V\mid(u,v)\in E\right\}\right|$ and $\operatorname{outdeg}_{G}(v)\coloneqq\left|\left\{u\in V\mid(v,u)\in E\right\}\right|$ for each $v\in V$ as the in-degree respectively out-degree of $v$ in $G$.
Each DAG $G$ can be associated with a partial order $\preceq_{G}$ on the vertex set $V(G)$, defined by $v\preceq_{G}w$ if and only if there is a directed $wv$-path. In this case, we say that $w$ is an ancestor of $v$ and $v$ is a descendant of $w$. If $v\preceq_{G}w$ and $v\neq w$, we write $v\prec_{G}w$. If $(u,v)\in E(G)$, then $u$ is a parent of $v$ and $v$ a child of $u$. We let $\operatorname{parent}_{G}(v)$ denote the set of all parents of $v$ in $G$. Two vertices $u,v\in V(G)$ are $\preceq_{G}$-incomparable if neither $u\preceq_{G}v$ nor $v\preceq_{G}u$ is true. We denote by $L(G)\subseteq V(G)$ the $\preceq_{G}$-minimal vertices of $G$, and we call $x\in L(G)$ a leaf of $G$. If $L(G)=X$, then $G$ is a DAG on $X$. A vertex $v\in V^{0}(G)\coloneqq V(G)\setminus L(G)$ is called an inner vertex. A vertex $v\in V(G)$ of a DAG $G$ that is $\preceq_{G}$-maximal is called a root, and the set of roots of $G$ is denoted by $R(G)$. Note that $L(G)\neq\emptyset$ and $R(G)\neq\emptyset$ for all DAGs $G$ [30]. A (rooted) network $N$ is a DAG for which $|R(N)|=1$, i.e., $N$ has a unique root $\rho\in V(N)$.
An edge $e=(u,w)$ in a DAG $G$ is a shortcut if there is a vertex $v\in V(G)$ such that $w\prec_{G}v\prec_{G}u$, i.e., there is a directed $uw$-path that does not contain the edge $e$. A DAG without shortcuts is shortcut-free ($\operatorname{sf}$). We denote with $\operatorname{sf}(G)$ the DAG obtained from $G$ by removing all shortcuts.
For a DAG $G$ we define $\operatorname{rev}(G)\coloneqq(V(G),\{(v,u)\mid(u,v)\in E(G)\})$ as the reverse of $G$. Note that a directed graph $G$ is a DAG if and only if it admits a topological order $\lessdot_{G}$ defined on $V(G)$, i.e., the vertices of $G$ can be totally ordered such that $(x,y)\in E(G)$ implies that $x\lessdot_{G}y$ [26]. It is easy to see $\lessdot_{\operatorname{rev}(G)}$ defined by $x\lessdot_{\operatorname{rev}(G)}y$ if and only $y\lessdot_{G}x$ for all $x,y\in V(G)$ yields a topological order on $\operatorname{rev}(G)$. Thus, we obtain
Observation 2.2.
Let $G$ be a DAG. Then $\operatorname{rev}(G)$ is a DAG and $\operatorname{rev}(\operatorname{rev}(G))=G$.
Clustering and Descendant Systems
For every $v\in V(G)$ in a DAG $G$, the set of its descendant leaves
$\operatorname{\mathtt{C}}_{G}(v)\coloneqq\{x\in L(G)\mid x\preceq_{G}v\}$
is a cluster of $G$. We write $\mathfrak{C}_{G}\coloneqq\{\operatorname{\mathtt{C}}_{G}(v)\mid v\in V(G)\}$ for the set of all clusters in $G$. Note that $\mathfrak{C}_{G}$ is grounded for all DAGs $G$. However, if a DAG $G$ has several roots, $\mathfrak{C}_{G}$ may not be rooted and, therefore, not a clustering system. For later reference, we provide the following results which shows, among others, that the set of clusters $\mathfrak{C}_{G}$ is invariant under shortcut-removal from $G$.
Lemma 2.3 ([30, L. 2.5]).
Let $G=(V,E)$ be a DAG. Then, for all $u,v\in V$, it holds that $u\prec_{G}v$ if and only if $u\prec_{\operatorname{sf}(G)}v$ and, for all $v\in V$, it holds that $\operatorname{\mathtt{C}}_{G}(v)=\operatorname{\mathtt{C}}_{\operatorname{sf}(G)}(v)$. In particular, $\operatorname{sf}(G)$ is a shortcut-free DAG.
Lemma 2.4 ([39, L. 1]).
For all DAGs $G$ and all $u,v\in V(G)$ it holds that $u\preceq_{G}v$ implies $\operatorname{\mathtt{C}}_{G}(u)\subseteq\operatorname{\mathtt{C}}_{G}(v)$.
There are DAGs $G$ containing $\preceq_{G}$-incomparable vertices $u$ and $v$ for which $\operatorname{\mathtt{C}}_{G}(u)\subseteq\operatorname{\mathtt{C}}_{G}(v)$ is satisfied. For example, the vertices $b$ and $c$ of the DAG $G$ in Figure 1 are $\preceq_{G}$-incomparable but satisfy $\operatorname{\mathtt{C}}_{G}(b)=\operatorname{\mathtt{C}}_{G}(c)$. The following property ensures that there is a directed path between two vertices $v$ and $u$ precisely if their clusters are comparable with respect to inclusion.
Definition 2.5 ([22]).
A DAG $G$ has the path-cluster-comparability (PCC) property if it satisfies, for all $v\in V(G): uu$ and $v$ are $\preceq_{G}$-comparable if and only if $\operatorname{\mathtt{C}}_{G}(u)\subseteq\operatorname{\mathtt{C}}_{G}(v)$ or $\operatorname{\mathtt{C}}_{G}(v)\subseteq\operatorname{\mathtt{C}}_{G}(u)$.
For a DAG $G$ and a vertex $v\in V(G)$, let
$\mathtt{D}_{G}(v)\coloneqq\{u\in V(G)\mid u\preceq_{G}v\}$
be the set of all descendants of $v$ in $G$. Moreover, the descendant system $\mathfrak{D}_{G}\coloneqq\{\mathtt{D}_{G}(v)\mid v\in V(G)\}$ consists of all descendant sets in $G$. In general $\mathfrak{D}_{G}$ is not grounded as, for all non-leaf vertices $v\in V(G)\setminus L(G)$, we have $\{v\}\notin\mathfrak{D}_{G}$ since $\{v\}\neq\mathtt{D}_{G}(u)$ for any $u\in V(G)$. Note also that for every vertex $v\in V(G)$ it holds that $\operatorname{\mathtt{C}}_{G}(v)\subseteq\mathtt{D}_{G}(v)$, with equality if and only if $v$ is a leaf. We note in passing that a construction resembling $\mathfrak{D}_{G}$ appears in [20, Sec. 6], albeit in a quite different context. Moreover, the descendant set $\mathfrak{D}_{G}$ of $G$ coincides with the collection of all so-called principal lower sets of the poset $(V(G),\preceq_{G})$, see e.g. [15]. The latter has, to the best of our knowledge, so far not been studied in the context of DAGs and least common ancestors. To keep this contribution self-consistent, we include proofs of some of the properties of $\mathfrak{D}_{G}$ that also appear as Exercise 1.16(ii) of [15].
Lemma 2.6.
Let $G$ be a DAG and $u,v\in V(G)$. Then, $\mathtt{D}_{G}(u)\neq\mathtt{D}_{G}(v)$ if and only if $u\neq v$. Moreover, $u\preceq_{G}v$ if and only if $\mathtt{D}_{G}(u)\subseteq\mathtt{D}_{G}(v)$. In particular, $\mathtt{D}_{G}(u)\subsetneq\mathtt{D}_{G}(v)$ implies that $v\notin\mathtt{D}_{G}(u)$. Consequently, the poset $(\mathfrak{D}_{G},\subseteq)$ is isomorphic to the poset $(V(G),\preceq_{G})$.
Proof.
Let $G$ be a DAG and $u,v\in V(G)$. Clearly, $\mathtt{D}_{G}(u)\neq\mathtt{D}_{G}(v)$ implies $u\neq v$. Assume that $u\neq v$. One of these cases must occur: (i) $u\prec_{G}v$, (ii) $v\prec_{G}u$ or (iii) $u$ and $v$ are $\preceq_{G}$-incomparable. Since $G$ is DAG, in Case (i) we have $v\not\preceq_{G}u$ and in Case (ii) we have $u\not\preceq_{G}v$. Thus, in Case (i), we have $v\in\mathtt{D}_{G}(v)$ but $v\notin\mathtt{D}_{G}(u)$ and, in Case (ii), we have $u\in\mathtt{D}_{G}(u)$ but $u\notin\mathtt{D}_{G}(v)$. In Case (iii) we have $u\in\mathtt{D}_{G}(u)$ but $u\notin\mathtt{D}_{G}(v)$. Hence, in all cases $\mathtt{D}_{G}(v)\neq\mathtt{D}_{G}(u)$.
Assume now that $u\preceq_{G}v$. In this case, all $w\preceq_{G}u$ trivially satisfy $w\preceq_{G}v$ and $\mathtt{D}_{G}(u)\subseteq\mathtt{D}_{G}(v)$ follows. Suppose now that $\mathtt{D}_{G}(u)\subseteq\mathtt{D}_{G}(v)$. Since, by definition, $u\in\mathtt{D}_{G}(u)$ it follows that $u\in\mathtt{D}_{G}(v)$. Consequently, $u\preceq_{G}v$ must hold. Finally, assume that $\mathtt{D}_{G}(u)\subsetneq\mathtt{D}_{G}(v)$. Since $u\in\mathtt{D}_{G}(u)$, we have $u\in\mathtt{D}_{G}(v)$. Hence, $u\prec_{G}v$. This together with the fact that $G$ is a DAG implies that $v\not\prec_{G}u$ and thus, $v\notin\mathtt{D}_{G}(u)$.
Finally, the previous arguments also show that the map $\varphi:V(G)\to\mathfrak{D}_{G}$ defined by $v\mapsto\mathtt{D}_{G}(v)$ is a bijection that satisfy $u\preceq_{G}v$ if and only if $\mathtt{D}_{G}(u)\subseteq\mathtt{D}_{G}(v)$. In other words, $\varphi$ is an order-preserving bijection and thus an isomorphism of the posets $(V(G),\preceq_{G})$ and $(\mathfrak{D}_{G},\subseteq)$. ∎
There is a close connection between the descendant system $\mathfrak{D}_{G}$ and the set of clusters $\mathfrak{C}_{G}$. Note that $\mathfrak{D}_{G}$ is a set system on $V(G)$, while $\mathfrak{C}_{G}$ is a set system on $L(G)\subseteq V(G)$. In particular, $x\in L(G)$ if and only if $\{x\}=\mathtt{D}_{G}(x)\in\mathfrak{D}_{G}$. Moreover, one easily verifies that $\operatorname{\mathtt{C}}_{G}(v)=\mathtt{D}_{G}(v)\cap L(G)$ for every $v\in V(G)$ and, therefore, that $C\in\mathfrak{C}_{G}$ if and only if $C=D\cap L(G)$ for some $D\in\mathfrak{D}_{G}$. Therefore, we obtain
Observation 2.7.
For every DAG $G$ it holds that $\mathfrak{C}_{G}=\{D\cap L(G)\,:\,D\in\mathfrak{D}_{G}\}$.
By Observation 2.7, it is straightforward of obtain the set of clusters $\mathfrak{C}_{G}$ from the descendant system $\mathfrak{D}_{G}$. As we shall see later in Section 7, it is also possible to construct $\mathfrak{D}_{G}$ from the set of clusters of so-called leaf-extended versions $\operatorname{lxt}(G)$ of a DAG $G$, although the procedure is less simple (cf. Lemma 7.1).
The descendant system of a DAG $G$ contains much more information than the corresponding set of clusters of $G$. Still, descendant systems neglect some of the structure of $G$. In particular, we have $\mathfrak{D}_{G}=\mathfrak{D}_{G^{\prime}}$ whenever $G^{\prime}$ is obtained from $G$ by adding shortcuts (cf. Lemma 2.3). Nevertheless, every DAG $G$ is uniquely determined by $\mathfrak{D}_{G}$ up to its shortcuts, as shown next.
Lemma 2.8.
For a DAG $G$ we have $V(G)=\cup_{D\in\mathfrak{D}_{G}}D$. Let $E^{\prime}\subseteq E(G)$ denote the set of all edges in $G$ that are not shortcuts. Then, for any two vertices $u,v\in V(G)$, it holds that $(u,v)\in E^{\prime}$ if and only if $v\in\mathtt{D}_{G}(u)$, and for all $w\in\mathtt{D}_{G}(u)\setminus\{u,v\}$, we have $v\notin\mathtt{D}_{G}(w)$.
Proof.
Let $G$ be a DAG. Since $v\in\mathtt{D}_{G}(v)$ and $\mathtt{D}_{G}(v)\in\mathfrak{D}_{G}$ for all $v\in V(G)$ it follows that $V(G)=\cup_{v\in V(G)}\mathtt{D}_{G}(v)=\cup_{D\in\mathfrak{D}_{G}}D$. Let $E^{\prime}\subseteq E(G)$ denote the set of all edges in $G$ that are not shortcuts. Let $(u,v)\in E^{\prime}$. Clearly, $v\in\mathtt{D}_{G}(u)$ holds. Let $w\in\mathtt{D}_{G}(u)$ with $w\notin\{u,v\}$. Since $(u,v)$ is not a shortcut in $G$, it cannot hold that $v\prec_{G}w\prec_{G}u$. Hence, $v\notin\mathtt{D}_{G}(w)$. Suppose now that $u,v\in V(G)$ are distinct vertices for which it holds that $v\in\mathtt{D}_{G}(u)$ and for all $w\in\mathtt{D}_{G}(u)\setminus\{u,v\}$ we have $v\notin\mathtt{D}_{G}(w)$. Hence, $v\prec_{G}u$ and no descendant $w\notin\{u,v\}$ of $u$ is an ancestor of $v$ in $G$. Hence, there is no vertex $w\in V(G)$ that satisfies $v\prec_{G}w\prec_{G}u$. This and $v\prec_{G}u$ now implies that $(u,v)\in E(G)$ and, in particular, that $(u,v)$ is not a shortcut. ∎
Least Common Ancestors (LCAs)
For a given DAG $G$ and a non-empty subset $A\subseteq V(G)$, a vertex $v\in V(G)$ is a common ancestor of $A$ if $v$ is an ancestor of every vertex in $A$. We denote with $\mathtt{ANC}_{G}(A)$ the set of all common ancestors of $A$ in $G$. Moreover, $v$ is a least common ancestor (LCA) of $A$ if $v$ is a $\preceq_{G}$-minimal common ancestor of $A$. The set $\operatorname{LCA}_{G}(A)$ comprises all LCAs of $A$ in $G$. In a network $N$, the unique root is a common ancestor for all non-empty $A\subseteq V(N)$ and, therefore, $\operatorname{LCA}_{N}(A)\neq\emptyset$ for all non-empty $A\subseteq V(N)$. In this contribution, we will encounter several situations in which there is a unique LCA of certain subsets of the vertices, i.e. when $|\operatorname{LCA}_{G}(A)|=1$ for some non-empty subset $A\subseteq V(G)$. For simplicity, we will write $\operatorname{lca}_{G}(A)=v$ in case that $\operatorname{LCA}_{G}(A)=\{v\}$ and say that $\operatorname{lca}_{G}(A)$ is well-defined; otherwise, we leave $\operatorname{lca}_{G}(A)$ undefined. Moreover, we will write $\operatorname{lca}_{G}(u,v)$ instead of $\operatorname{lca}_{G}(\{u,v\})$ for $u,v\in V(G)$.
The following definition specifies DAGs where the LCA is unique for all non-empty (size-2) subsets of the leaf set. They have been explicitly studied in [22, 21, 39].
Definition 2.9.
Let $G=(V,E)$ be a DAG on $X$. Then,
  • $G$ has the lca-property if $|\operatorname{LCA}_{G}(A)|=1$ for all non-empty subsets $A\subseteq X$.
  • $G$ has pairwise-lca-property if $|\operatorname{LCA}_{G}(A)|=1$ for all subsets $A\subseteq X$ of size $|A|=2$.
Networks that have the lca-property are also called lca-networks. Moreover, $G$ is lca-relevant if, for all $v\in V(G)$, there is some subset $A\subseteq L(G)$ with $v=\operatorname{lca}_{G}(A)$. We emphasize that a DAG can be lca-relevant without having the lca-property, and vice versa. For examples and details on the relationship between these two properties we refer to [30, 21]. The following result generalizes Lemma 39 in [22], extending the result from networks to general DAGs.
Lemma 2.10.
If a DAG $G$ has the lca-property, then $\mathfrak{C}_{G}$ is a closed clustering system.
Proof.
Let $G$ be a DAG on $X$ with the lca-property. Clearly, $\emptyset\notin\mathfrak{C}_{G}$ and $\{x\}\in\mathfrak{C}_{G}$ for each $x\in X$ and, thus, $\mathfrak{C}_{G}$ is grounded. To show that $\mathfrak{C}_{G}$ is a clustering system it thus suffices to show that $X\in\mathfrak{C}_{G}$, i.e., that $\mathfrak{C}_{G}$ is rooted. Since $G$ has the lca-property, $v=\operatorname{lca}_{G}(X)$ must be well-defined. By definition of least common ancestors, $X\subseteq\operatorname{\mathtt{C}}_{G}(v)$ which, together with $\operatorname{\mathtt{C}}_{G}(v)\subseteq X$, implies that $\operatorname{\mathtt{C}}_{G}(v)=X$. Hence, $X\in\mathfrak{C}_{G}$. In summary, $\mathfrak{C}_{G}$ is a clustering system.
It remains to show that $\mathfrak{C}_{G}$ is closed. Let $\operatorname{\mathtt{C}}_{G}(u),\operatorname{\mathtt{C}}_{G}(v)\in\mathfrak{C}_{G}$ for some $u,v\in V(G)$. If $\operatorname{\mathtt{C}}_{G}(u)\subseteq\operatorname{\mathtt{C}}_{G}(v)$, $\operatorname{\mathtt{C}}_{G}(v)\subseteq\operatorname{\mathtt{C}}_{G}(u)$ or $\operatorname{\mathtt{C}}_{G}(u)\cap\operatorname{\mathtt{C}}_{G}(v)=\emptyset$, there is nothing to show. Hence, assume that $\operatorname{\mathtt{C}}_{G}(u)\cap\operatorname{\mathtt{C}}_{G}(v)\notin\{\operatorname{\mathtt{C}}_{G}(u),\operatorname{\mathtt{C}}_{G}(v),\emptyset\}$ and put $A\coloneqq\operatorname{\mathtt{C}}_{G}(u)\cap\operatorname{\mathtt{C}}_{G}(v)\neq\emptyset$. Since $G$ has the lca-property, there is $w\in V(G)$ such that $w=\operatorname{lca}_{G}(A)$, and thus $A\subseteq\operatorname{\mathtt{C}}_{G}(w)$. The contraposition of Lemma 2.4 shows that $u$ and $v$ are two $\preceq_{G}$-incomparable common ancestors of $A$. Since $w$ is the unique $\preceq_{G}$-minimal common ancestor of $A$, we have $w\preceq_{G}u$ and $w\preceq_{G}v$, which together with Lemma 2.4 implies $\operatorname{\mathtt{C}}_{G}(w)\subseteq\operatorname{\mathtt{C}}_{G}(u)$ and $\operatorname{\mathtt{C}}_{G}(w)\subseteq\operatorname{\mathtt{C}}_{G}(v)$. Therefore $\operatorname{\mathtt{C}}_{G}(w)\subseteq A$. Hence $A=\operatorname{\mathtt{C}}_{G}(w)\in\mathfrak{C}_{G}$ and thus, $\mathfrak{C}_{G}$ is closed. ∎
The converse of Lemma 2.10 is not true. A counter-example is shown in Figure 1.
Figure 1: The clustering system $\mathfrak{C}_{G}=\{\{x\},\{y\},\{x,y\}\}$, indicated in blue, of $G$ is closed but $G$ does not have the lca-property as $|\operatorname{LCA}_{G}(\{x,y\})|>1$. Note that $G$ contains a $K_{2,2}$-minor without $\mathsf{X}$}- or $\mathsf{X^{\prime}}$}-subdivision, see Section 4 for more details. Although $\mathfrak{C}_{G}$ is closed, the descendant system $\mathfrak{D}_{G}$ of $G$ is not closed because $\mathtt{D}_{G}(b)\cap\mathtt{D}_{G}(c)=\{b,x,y\}\cap\{c,x,y\}=\{x,y\}$ but $\{x,y\}\notin\mathfrak{D}_{G}$.
The next rather simple but technical result will be helpful to simplify several arguments throughout this paper.
Lemma 2.11.
Let $G$ be a DAG such that, for some $x,y\in V(G)$, $\operatorname{lca}_{G}(x,y)$ is well-defined. If $u,v\in V(G)$ are $\preceq_{G}$-incomparable vertices of $G$ such that both $u$ and $v$ are common ancestors of $\{x,y\}$, then $\operatorname{lca}_{G}(x,y)\prec_{G}u$ and $\operatorname{lca}_{G}(x,y)\prec_{G}v$.
Proof.
By definition of least common ancestors, $\operatorname{lca}_{G}(x,y)\preceq_{G}v$ and $\operatorname{lca}_{G}(x,y)\preceq_{G}u$ holds. Therefore, if $\operatorname{lca}_{G}(x,y)\in\{u,v\}$, then $u$ and $v$ are $\preceq_{G}$-comparable. Since $u$ and $v$ are $\preceq_{G}$-incomparable, it thus follows that $\operatorname{lca}_{G}(x,y)\prec_{G}v$ and $\operatorname{lca}_{G}(x,y)\prec_{G}u$. ∎
For a poset $(\mathfrak{Q},\sqsubseteq)$, the Hasse diagram $\mathscr{H}(\mathfrak{Q},\sqsubseteq)$ is the DAG with vertex set $\mathfrak{Q}$ and directed edges from $A\in\mathfrak{Q}$ to $B\in\mathfrak{Q}$ if (i) $B\sqsubseteq A$ and $A\neq B$ and (ii) there is no $C\in\mathfrak{Q}$ with $B\sqsubseteq C\sqsubseteq A$ and $C\neq A,B$.
A DAG $G=(V,E)$ is regular if the map $\varphi\colon V\to V(\mathscr{H}(\mathfrak{C}_{G},\subseteq))$ defined by $v\mapsto\operatorname{\mathtt{C}}_{G}(v)$ is an isomorphism between $G$ and $\mathscr{H}(\mathfrak{C}_{G},\subseteq)$ where $\subseteq$ denotes the usual subset-relation. The next result summarizes some of the results from [30, Sec. 4].
Theorem 2.12.
The following statements are equivalent for every DAG $G$.
  1. (1)
    $G$ is lca-relevant
  2. (2)
    $v=\operatorname{lca}_{G}(\operatorname{\mathtt{C}}_{G}(v))$ for all $v\in V(G)$ (known as the strong-(CL) property)
  3. (3)
    $G$ satisfies (PCC) and $\operatorname{\mathtt{C}}_{G}(u)\neq\operatorname{\mathtt{C}}_{G}(v)$ for all distinct $u,v\in V(G)$.
  4. (4)
    $\operatorname{\mathtt{C}}_{G}(u)\subseteq\operatorname{\mathtt{C}}_{G}(v)$ if and only if $u\preceq_{G}v$ for all $u,v\in V(G)$.
  5. (5)
    $G$ is isomorphic to the regular DAG $\mathscr{H}(\mathfrak{C}_{G},\subseteq)$ to which $k\geq 0$ shortcuts have been added.
3 Basic properties of global lca-networks and a first characterization
The lca-property can be strengthened as follows.
Definition 3.1.
Let $G=(V,E)$ be a DAG on $X$. Then,
  • $G$ has the global lca-property if $|\operatorname{LCA}_{G}(A)|=1$ for all non-empty subsets $A\subseteq V$.
  • $G$ has the global pairwise-lca-property if $|\operatorname{LCA}_{G}(A)|=1$ for all subsets $A\subseteq V$ of size $|A|=2$.
Moreover, DAGs that have the global lca-property are also called global lca-DAGs.
The next result will be used throughout the paper without explicit mention each time.
Observation 3.2.
Every global lca-network is an lca-network.
Even more interesting is the following
Lemma 3.3.
If a DAG $G$ on $X$ has the global lca-property then $G$ is a network and $\mathfrak{C}_{G}$ is a closed clustering system on $X$.
Proof.
Let $G$ be a DAG on $X$ with global lca-property. Suppose that $G$ is not a network and thus, $|R(G)|>1$. By definition, $\operatorname{LCA}_{G}(R(G))=\emptyset$. Hence, $G$ does not have the global lca-property; a contradiction. Hence, $G$ must be a network. By [22, L. 14], for any network $G$, the set system $\mathfrak{C}_{G}$ is a clustering system. Moreover, since $G$ has, in particular, the lca-property, Lemma 2.10 implies that $\mathfrak{C}_{G}$ is closed. ∎
We note that, similar to the ideas in [30], we may define the global LCA-property as a property of a DAG $G$ to ensure that for all subsets (possibly of a given size) we have $\operatorname{LCA}(A)\neq\emptyset$ for all $A\subseteq V(G)$, i.e., a not necessarily unique LCA for all subsets $A\subseteq V(G)$ exists. However, by similar arguments that lead to Lemma 3.3, each DAG $G$ with the global LCA-property must be a network with a unique root $\rho_{G}$. In this case, however, $\operatorname{\mathtt{C}}_{G}(\rho_{G})=L(G)$, which implies that $\rho_{G}$ serves as an ancestor of all subsets $A$. Consequently, $\operatorname{LCA}(A)\neq\emptyset$ for all $A\subseteq V(G)$ is trivially satisfied if and only if $G$ is a network. Thus, we focus here on the more interesting case of the global lca-property as in Definition 3.1.
Remark 1.
Since global lca-DAGs must be networks, we call such DAGs also global lca-networks.
The converse of Lemma 3.3 is not true. In particular, Figure 1 shows a DAG $G$, for which $\mathfrak{C}_{G}$ is a closed clustering system but where $G$ does not have the lca-property and thus, in particular, not the global lca-property. As we shall see later in Section 7, closedness of $\mathfrak{D}_{G}$ instead characterizes global lca-networks. However, in case that $G$ satisfies (PCC), closedness of its clustering system characterizes global lca-network as shown in the next result, which generalizes Theorem 7 in [22].
Proposition 3.4.
Let $G$ be a DAG satisfying (PCC). If $\mathfrak{C}_{G}$ is a clustering system, then $G$ is a network. Moreover, $G$ is a global lca-network if and only if $\mathfrak{C}_{G}$ is a closed clustering system.
Proof.
Let $G$ be a DAG satisfying (PCC). Suppose first that $\mathfrak{C}_{G}$ is a clustering system. Let $v\in V(G)$ be a vertex such that $\operatorname{\mathtt{C}}_{G}(v)=X$, which exists since $\mathfrak{C}_{G}$ is rooted. Clearly, there is some root $r\in R(G)$ such that $v\preceq_{G}r$. By Lemma 2.4, this vertex $r$ must satisfy $X=\operatorname{\mathtt{C}}_{G}(v)\subseteq\operatorname{\mathtt{C}}_{G}(r)\subseteq X$ and thus, $\operatorname{\mathtt{C}}_{G}(r)=X$. Since $r$ is a root in $G$, there is no vertex $u$ satisfying $r\prec_{G}u$. This together with (PCC) and the fact that $\operatorname{\mathtt{C}}_{G}(u)\subseteq X$ for all vertices $u$ in $G$ implies that $u\preceq_{G}r$ for all vertices $u$ in $G$. Hence $r$ is the unique root in $G$ and $G$ is a network.
Suppose now that $\mathfrak{C}_{G}$ is a closed clustering system. By the latter arguments, $G$ is a network and thus, $\operatorname{LCA}_{G}(A)\neq\emptyset$ for all $A\subseteq V(G)$. Assume, for contradiction, that there is a non-empty set $A\subseteq V(G)$ with $|\operatorname{LCA}_{G}(A)|>1$. Let $u,v\in\operatorname{LCA}_{G}(A)$ be two distinct vertices. Then $u,v$ must be $\preceq_{G}$-incomparable. Let $a\in A$. Since $u,v\in\operatorname{LCA}_{G}(A)$, it follows that $a\preceq_{G}u$ and $a\preceq_{G}v$. Lemma 2.4 implies that $\operatorname{\mathtt{C}}_{G}(a)\subseteq\operatorname{\mathtt{C}}_{G}(u)$ and $\operatorname{\mathtt{C}}_{G}(a)\subseteq\operatorname{\mathtt{C}}_{G}(v)$. Hence, $C\coloneqq\operatorname{\mathtt{C}}_{G}(u)\cap\operatorname{\mathtt{C}}_{G}(v)\neq\emptyset$. Since $\mathfrak{C}_{G}$ is closed, $C\in\mathfrak{C}_{G}$. Thus, there is a vertex $w\in V(G)$ such that $\operatorname{\mathtt{C}}_{G}(w)=C$. Note that, since $G$ satisfies (PCC), all vertices $w^{\prime},w^{\prime\prime}$ with $\operatorname{\mathtt{C}}_{G}(w^{\prime})=\operatorname{\mathtt{C}}_{G}(w^{\prime\prime})=C$ must be $\preceq_{G}$-comparable. Thus, we can assume without loss of generality that $w$ is the unique $\preceq_{G}$-maximal vertex satisfying $\operatorname{\mathtt{C}}_{G}(w)=C$. Note that neither $\operatorname{\mathtt{C}}_{G}(u)\subseteq\operatorname{\mathtt{C}}_{G}(v)$ nor $\operatorname{\mathtt{C}}_{G}(v)\subseteq\operatorname{\mathtt{C}}_{G}(u)$ is possible because $G$ satisfies (PCC) and $u,v$ are $\preceq_{G}$-incomparable. This and $\operatorname{\mathtt{C}}_{G}(u)\cap\operatorname{\mathtt{C}}_{G}(v)=\operatorname{\mathtt{C}}_{G}(w)$ implies that neither $\operatorname{\mathtt{C}}_{G}(u)=\operatorname{\mathtt{C}}_{G}(w)$ nor $\operatorname{\mathtt{C}}_{G}(v)=\operatorname{\mathtt{C}}_{G}(w)$ can hold. Hence, $\operatorname{\mathtt{C}}_{G}(w)\subsetneq\operatorname{\mathtt{C}}_{G}(u)$ and $\operatorname{\mathtt{C}}_{G}(w)\subsetneq\operatorname{\mathtt{C}}_{G}(v)$. In particular, $w\neq u$ and $w\neq v$. The latter two arguments together with (PCC) and Lemma 2.4 implies that $w\prec_{G}u$ and $w\prec_{G}v$. Now let $b\in A$, where $b=a$ may be possible. Since $u,v\in\operatorname{LCA}_{G}(A)$, it follows that $b\preceq_{G}u$ and $b\preceq_{G}v$. Lemma 2.4 implies that $\operatorname{\mathtt{C}}_{G}(b)\subseteq\operatorname{\mathtt{C}}_{G}(u)$ and $\operatorname{\mathtt{C}}_{G}(b)\subseteq\operatorname{\mathtt{C}}_{G}(v)$. Therefore, $\operatorname{\mathtt{C}}_{G}(b)\subseteq\operatorname{\mathtt{C}}_{G}(u)\cap\operatorname{\mathtt{C}}_{G}(v)=\operatorname{\mathtt{C}}_{G}(w)$. Since $G$ satisfies (PCC), $b$ and $w$ must be $\preceq_{G}$-comparable. If $w\prec_{G}b$, then Lemma 2.4 implies that $\operatorname{\mathtt{C}}_{G}(w)\subseteq\operatorname{\mathtt{C}}_{G}(b)$ and thus, $\operatorname{\mathtt{C}}_{G}(w)=\operatorname{\mathtt{C}}_{G}(b)$; a contradiction to $w$ being the $\preceq_{G}$-maximal vertex satisfying $\operatorname{\mathtt{C}}_{G}(w)=C$. Hence, $b\preceq_{G}w$ must hold. As $b\in A$ was chosen arbitrarily, it follows that for all $b\in A$ it holds that $b\preceq_{G}w$. Hence, $w$ is a common ancestor of the elements in $A$. However, since $w\prec_{G}u$ and $w\prec_{G}v$, we have $u,v\notin\operatorname{LCA}_{G}(A)$; a contradiction. In summary, $|\operatorname{LCA}_{G}(A)|=1$ for all non-empty subsets $A\subseteq V(G)$. Therefore, $G$ is a global lca-network. Finally, if $G$ is a global lca-network, then Lemma 3.3 implies that $\mathfrak{C}_{G}$ is a closed clustering system. ∎
We now provide a first characterization of global lca-networks. To this end, we will adapt ideas established for semilattices (cf. e.g. [17, page 4]).
Theorem 3.5.
The following statements are equivalent for every DAG $G$.
  1. (1)
    $G$ has the global pairwise-lca-property
  2. (2)
    $G$ has the global lca-property.
Proof.
Clearly, Condition (2) implies (1). We show now that Condition (1) implies (2). To this end, suppose that $G$ is a DAG that has the global pairwise-lca-property. Clearly, $\operatorname{lca}_{G}(\{x\})=x$ holds for all $x\in V(G)$. Consider now $W=\{a_{1},a_{2},a_{3}\}\subseteq V(G)$. Hence, $\operatorname{lca}_{G}(a_{1},a_{2})=v$ and $\operatorname{lca}_{G}(a_{3},v)=w$ exist. We claim that $w=\operatorname{lca}_{G}(\{a_{1},a_{2},a_{3}\})$. To see this, first observe that, by choice of $v$ and $w$, we have $a_{1},a_{2}\preceq_{G}v$ respectively $v,a_{3}\preceq_{G}w$ which further implies that $a_{1},a_{2},a_{3}\preceq_{G}w$ and thus, $w$ is a common ancestor of $W$. Now, let $u$ be an arbitrary common ancestor of $a_{1},a_{2}$ and $a_{3}$. Thus, $a_{1},a_{2},a_{3}\preceq_{G}u$ and therefore $u$ is, in particular, a common ancestor of $a_{1}$ and $a_{2}$. By definition of least common ancestors, $\operatorname{lca}_{G}(a_{1},a_{2})=v\preceq_{G}u$ holds. Since $v\preceq_{G}u$ and $a_{3}\preceq_{G}u$, the definition of least common ancestors furthermore implies that $w=\operatorname{lca}_{G}(a_{3},v)\preceq_{G}u$. As these arguments hold for any common ancestor $u$ of $a_{1},a_{2},a_{3}$, we can conclude that $w=\operatorname{lca}_{G}(a_{3},v)=\operatorname{lca}_{G}(a_{1},a_{2},a_{3})$. Applying the latter arguments inductively, one arrives at $\operatorname{lca}_{G}(W)=\operatorname{lca}_{G}(\dots(\operatorname{lca}_{G}(a_{1},a_{2}),a_{3}),\dots,a_{\ell})$ for all non-empty sets $W=\{a_{1},\dots,a_{\ell}\}\subseteq V(G)$. Hence, $G$ has the global lca-property. ∎
As a consequence of Theorem 3.5, it suffices to check the existence of a unique least common ancestor for all pairs of vertices to conclude that a DAG has the global lca-property. Since Kowaluk and Lingas [28, Thm. 3] have shown that verifying the existence of a unique LCA for all pairs of vertices of a given DAG can be achieved in polynomial time, we immediately obtain the following result.
Corollary 3.6.
It can be verified in polynomial time if a DAG $G$ has the global lca-property.
The remaining three results of this section formalize the, intuitively clear, connections between join-semilattices and global lca-networks:
Proposition 3.7.
For every poset $(\mathfrak{Q},\sqsubseteq)$ the following statements are equivalent.
  1. (1)
    The Hasse diagram $\mathscr{H}(\mathfrak{Q},\sqsubseteq)$ has the global pairwise-lca-property.
  2. (2)
    The Hasse diagram $\mathscr{H}(\mathfrak{Q},\sqsubseteq)$ has the global lca-property.
  3. (3)
    $(\mathfrak{Q},\sqsubseteq)$ is a join-semilattice.
Proof.
By Theorem 3.5, Statement (1) and (2) are equivalent. Suppose that the Hasse diagram $\mathscr{H}\coloneqq\mathscr{H}(\mathfrak{Q},\sqsubseteq)$ has the global pairwise-lca-property. Note that $X\preceq_{\mathscr{H}}Y$ if and only if $X\sqsubseteq Y$ for all $X,Y\in\mathfrak{Q}$. Moreover, since the Hasse diagram $\mathscr{H}$ has the global pairwise-lca-property, for all $X,Y\in\mathfrak{Q}$, $Z_{XY}\coloneqq\operatorname{lca}_{\mathscr{H}}(X,Y)$ is well-defined. By definition, $Z_{XY}=\sup_{\mathfrak{Q}}(X,Y)$ and thus, $(\mathfrak{Q},\sqsubseteq)$ is a join-semilattice. Therefore, (1) implies (3). Suppose that (3) holds. Thus, for $X_{1},X_{2}\in\mathfrak{Q}$ there is a unique least upper bound $\sup_{\mathfrak{Q}}(X_{1},X_{2})\in\mathfrak{Q}$. It now immediately follows that $\sup_{\mathfrak{Q}}(X_{1},X_{2})$ is a vertex in $\mathscr{H}$. Moreover, by definition of the Hasse diagram, $\operatorname{lca}_{\mathscr{H}}(X_{1},X_{2})=\sup_{\mathfrak{Q}}(X_{1},X_{2})$. Hence, $\mathscr{H}$ has the global pairwise-lca-property. ∎
Corollary 3.8.
If a set system $\mathfrak{Q}$ is grounded and $(\mathfrak{Q},\subseteq)$ is a join-semilattice, then $\mathfrak{Q}$ is a clustering system and there is a DAG $G$ with $\mathfrak{C}_{G}=\mathfrak{Q}$ and with global pairwise-lca-property. In this case, $G$ is a network.
Proof.
Assume that $\mathfrak{Q}\subseteq 2^{X}$ is a grounded set system on $X$ and $(\mathfrak{Q},\subseteq)$ is a join-semilattice. Since $\mathfrak{Q}$ is grounded, we have $\{x\}\in\mathfrak{Q}$ for all $x\in X$. Consider the graph $G\coloneqq\mathscr{H}(\mathfrak{Q},\subseteq)$, obtained by replacing each singleton vertex $\{x\}$ in $\mathscr{H}(\mathfrak{Q},\subseteq)$ by the element $x$. Since $L(\mathscr{H}(\mathfrak{Q},\subseteq))=\{\{x\}\mid x\in X\}$ follows from $\mathfrak{Q}$ being grounded (in particular, $\emptyset\notin\mathfrak{Q}$) we have, together with the definition of Hasse diagrams, $L(G)=X$. Moreover, the construction of $G$ ensures that $\mathfrak{C}_{G}=\mathfrak{Q}$. Since $(\mathfrak{Q},\subseteq)$ is a join-semilattice, Proposition 3.7 implies that $\mathscr{H}(\mathfrak{Q},\subseteq)$ has the global pairwise-lca-property. Since $G$ and $\mathscr{H}(\mathfrak{Q},\subseteq)$ are isomorphic, $G$ has the global pairwise-lca-property. By Lemma 3.3, $G$ is a network and $\mathfrak{C}_{G}=\mathfrak{Q}$ is a clustering system. ∎
Theorem 3.9.
A network $N$ has the global lca-property if and only if $(V(N),\preceq_{N})$ is a join-semilattice.
Proof.
One easily verifies that $\operatorname{sf}(N)=\mathscr{H}((V(N),\preceq_{N}))$. By Lemma 2.3, NN has the global lca-property if and only if $\operatorname{sf}(N)=\mathscr{H}((V(N),\preceq_{N}))$ has the global lca-property. By Proposition 3.7, the latter is equivalent to $(V(N),\preceq_{N})$ being a join-semilattice. ∎
4 A characterization by $K_{2,2}$-minors
An edge-subdivision of a graph $G$ is the replacement of an edge $(x,y)$ in $G$ by the two edges $(x,z)$ and $(z,y)$ for some new vertex $z\notin V(G)$. A graph $G^{\prime}$ is subdivision of $G$ if $G^{\prime}$ can be obtained from $G$ by a sequence of edge-subdivisions. A graph $H$ is a (topological) minor of $G$, also known as a homeomorphic subgraph, if a subdivision of $H$ is isomorphic to a subgraph $G^{\prime}$ of $G$. In this case, we say that $G$ contains an $H$-minor and that $G^{\prime}$ is an $H$-subdivision. A $K_{2,2}$ is the DAG having two roots $r_{1}$ and $r_{2}$, two leaves $\ell_{1}$ and $\ell_{2}$ and edges ($r_{i},\ell_{j})$, $1\leq i,j\leq 2$, depicted in Figure 2. At a first glance, one might assume that the appearance of a $K_{2,2}$-minor in a DAG $G$ indicates the existence of two vertices $x$ and $y$ such that $|\operatorname{LCA}_{G}(\{x,y\})|\geq 2$. However, the network $N$ of Figure 2 exemplifies that this is not the case; $N$ has a $K_{2,2}$-minor yet still has the global lca-property. We need the following additional structure on the $K_{2,2}$-minors to find a connection to unique least common ancestors.
Definition 4.1.
A DAG $G$ contains a strict $K_{2,2}$-minor if it contains $K_{2,2}$-subdivision $H\subseteq G$ such that $r$ and $r^{\prime}$, respectively $\ell$ and $\ell^{\prime}$, are $\preceq_{G}$-incomparable for $R(H)=\{r,r^{\prime}\}$ and $L(H)=\{\ell,\ell^{\prime}\}$. In this case, we call $H$ a strict $K_{2,2}$-subdivision.
Lemma 4.2.
Let $G=(V,E)$ be a DAG. If there exists $u,v\in V$ such that $|\operatorname{LCA}_{G}(\{u,v\})|\geq 2$, then $G$ contains a $K_{2,2}$-minor. In particular, there is a strict $K_{2,2}$-subdivision $H\subseteq G$ such that $R(H)\subseteq\operatorname{LCA}_{G}(\{u,v\})\cap\operatorname{LCA}_{G}(L(H))$.
Proof.
Let $G=(V,E)$ be a DAG and assume $u,v\in V$ are vertices such that $|\operatorname{LCA}_{G}(\{u,v\})|\geq 2$. In particular, $u\neq v$ holds and we may pick distinct $r,r^{\prime}\in\operatorname{LCA}_{G}(\{u,v\})$. Now, since $u,v\prec_{G}r$ and $u,v\prec_{G}r^{\prime}$ four directed paths necessarily exist in $G$: $P_{ru}$, $P_{rv}$, $P_{r^{\prime}u}$ and $P_{r^{\prime}v}$, where $P_{qw}$ denotes some path from $q\in\{r,r^{\prime}\}$ to $w\in\{u,v\}$. Since $r$ and $r^{\prime}$ are both LCAs of $\{u,v\}$, no vertex $p\prec_{G}r$ is a common ancestor of both $u$ and $v$, and no vertex $p^{\prime}\prec_{G}r^{\prime}$ is a common ancestor of both $u$ and $v$. It is not difficult to see that this implies that
$V(P_{ru})\cap V(P_{rv})=\{r\},\quad V(P_{ru})\cap V(P_{r^{\prime}v})=\emptyset,\quad V(P_{r^{\prime}u})\cap V(P_{r^{\prime}v})=\{r^{\prime}\}\text{, and}\quad V(P_{r^{\prime}u})\cap V(P_{rv})=\emptyset.$ (1)
However, it is possible that the paths $P_{ru}$ and $P_{r^{\prime}u}$, or $P_{rv}$ and $P_{r^{\prime}v}$, resp., have multiple vertices in common. Let $u^{\prime}$ be the $\preceq_{G}$-maximal vertex in $V(P_{ru})\cap V(P_{r^{\prime}u})$ and let $P_{ru^{\prime}}$ and $P_{r^{\prime}u^{\prime}}$ be the subpaths of $P_{ru}$ respectively $P_{r^{\prime}u}$ from $r$ respectively $r^{\prime}$ to $u^{\prime}$. Similarly, let $v^{\prime}$ be the $\preceq_{G}$-maximal vertex in $V(P_{rv})\cap V(P_{r^{\prime}v})$ and define $P_{rv^{\prime}}$ and $P_{r^{\prime}v^{\prime}}$ analogously as the respective subpaths from $r$ to $v^{\prime}$ and $r^{\prime}$ to $v^{\prime}$. By definition, $V(P_{ru^{\prime}})\cap V(P_{r^{\prime}u^{\prime}})=\{u^{\prime}\}$ and $V(P_{rv^{\prime}})\cap V(P_{r^{\prime}v^{\prime}})=\{v^{\prime}\}$ holds. This argument together with Equ. (1) shows that the only common vertices of $P_{ru^{\prime}}$, $P_{r^{\prime}u^{\prime}}$, $P_{rv^{\prime}}$ and $P_{r^{\prime}v^{\prime}}$ are $r,r^{\prime},u^{\prime}$ and $v^{\prime}$. Hence, the subgraph $H$ of $G$ that is composed of the respective vertices and edges of the latter four paths form a subdivision of a $K_{2,2}$, i.e., $G$ contains $K_{2,2}$-minor. In particular $R(H)=\{r,r^{\prime}\}\subseteq\operatorname{LCA}_{G}(\{u,v\})$ and $L(H)=\{u^{\prime},v^{\prime}\}$ holds by construction. By definition of LCAs, $r$ and $r^{\prime}$ must be $\preceq_{G}$-incomparable. Since both $u\preceq_{G}u^{\prime}\prec_{G}r$, $v\preceq_{G}v^{\prime}\prec_{G}r$ and $r\in\operatorname{LCA}_{G}(\{u,v\})$ holds, it follows that $r\in\operatorname{LCA}_{G}(\{u^{\prime},v^{\prime}\})$. Similarly, $r^{\prime}\in\operatorname{LCA}_{G}(\{u^{\prime},v^{\prime}\})$, hence $R(H)\subseteq\operatorname{LCA}_{G}(\{u,v\})\cap\operatorname{LCA}_{G}(L(H))$. In particular, $|\operatorname{LCA}_{G}(\{u^{\prime},v^{\prime}\})|\geq|\{r,r^{\prime}\}|=2$. Therefore, $u^{\prime}$ and $v^{\prime}$ must be $\preceq_{G}$-incomparable. In summary, $H$ is a subdivision of a $K_{2,2}$ such that the vertices in $R(H)$, resp., $L(H)$ are $\preceq_{G}$-incomparable. Hence, $G$ contains a strict $K_{2,2}$-minor. ∎
Note that the contrapositive of the first statement of Lemma 4.2 implies that if $N=(V,E)$ is a network that does not have a $K_{2,2}$ as a minor, then $|\operatorname{LCA}_{N}(\{u,v\})|\leq 1$ for all $\{u,v\}\in\binom{V}{2}$. Since $|\operatorname{LCA}_{N}(\{u,v\})|>0$ holds in all networks, we therefore have $|\operatorname{LCA}_{N}(\{u,v\})|=1$ for all $\{u,v\}\in\binom{V}{2}$ i.e $N$ has the global pairwise-lca-property. By Theorem 3.5, this implies that $\operatorname{lca}_{N}(Y)$ is well-defined for all non-empty $Y\subseteq V$. That is, we obtain
Corollary 4.3.
If a network $N$ does not have a $K_{2,2}$-minor, then $N$ is a global lca-network.
Neither Lemma 4.2 nor Corollary 4.3 can be strengthened to equivalences. To see this, consider the global lca-network $N^{\prime}$ in Figure 2, which contains a strict $K_{2,2}$-subdivision $H$ whose vertices and edges are indicated within gray-shaded area. In particular, the leaves $\ell$ and $\ell^{\prime}$ of $H$ satisfy $\operatorname{lca}_{N^{\prime}}(\ell,\ell^{\prime})=v$ and, informally speaking, the existence of this vertex $v$ is displayed as an $\mathsf{X}$-shaped pattern “inside” the $K_{2,2}$-minor. To make this statement more precise, we let $\mathsf{X}$ denote the DAG with five vertices $r_{1},r_{2},w,\ell_{1},\ell_{2}$ and directed edges $(r_{i},w)$ and $(w,\ell_{i})$, $1\leq i\leq 2$. Moreover, let $\mathsf{X}^{\prime}$ denote the DAG with six vertices $r_{1},r_{2},v,w,\ell_{1},\ell_{2}$ and directed edges $(v,w)$, $(r_{i},v)$ and $(w,\ell_{i})$, $1\leq i\leq 2$. See Figure 2 for depictions of the DAGs $\mathsf{X}$} and $\mathsf{X^{\prime}}$}.
Figure 2: The three DAGs $K_{2,2}$, $\mathsf{X}$ and $\mathsf{X}^{\prime}$ that appear in the characterization of global lca-networks in terms of minors (Thm. 4.4). In addition, two global lca-networks $N$ and $N^{\prime}$ are shown. Here, NN contains a $K_{2,2}$-minor but not a strict $K_{2,2}$-minor and $N^{\prime}$ contains a strict $K_{2,2}$-minor (the respective $K_{2,2}$-minors are located within the gray-shaded area).
Theorem 4.4.
A network $N$ has the global lca-property if and only if, for every strict $K_{2,2}$-subdivision $H\subseteq N$, there exists an $\mathsf{X}$- or $\mathsf{X^{\prime}}$-subdivision $H^{\prime}\subseteq N$ with $L(H)=L(H^{\prime})$ and $R(H)=R(H^{\prime})$.
Proof.
Let $N=(V,E)$ be a network. First assume that $N$ has the global lca-property. If $N$ does not contain any strict $K_{2,2}$-minor the statement is vacuously true. Hence assume that $N$ contains a strict $K_{2,2}$-minor. Let $H\subseteq N$ be a strict $K_{2,2}$-subdivision and let $R(H)=\{r,r^{\prime}\}$ and $L(H)=\{\ell,\ell^{\prime}\}$. By definition of strict $K_{2,2}$-minors, $r$ and $r^{\prime}$, resp., $\ell$ and $\ell^{\prime}$ are $\preceq_{N}$-incomparable. Since $N$ has the global lca-property, $w\coloneqq\operatorname{lca}_{N}(\ell,\ell^{\prime})$ is a well-defined vertex satisfying $\ell,\ell^{\prime}\prec_{N}w$, where $w\neq\ell,\ell^{\prime}$ follows from the fact that $\ell$ and $\ell^{\prime}$ are $\preceq_{N}$-incomparable. Hence two directed paths $P_{w\ell}$ and $P_{w\ell^{\prime}}$ exist in $N$, from $w$ to $\ell$ and $\ell^{\prime}$, respectively. Since $w$ is the least common ancestor of $\ell$ and $\ell^{\prime}$, it must hold that $V(P_{w\ell})\cap V(P_{w\ell^{\prime}})=\{w\}$. Since both $r$ and $r^{\prime}$ are $\preceq_{N}$-incomparable common ancestors of $\{\ell,\ell^{\prime}\}$, Lemma 2.11 implies that $w\prec_{N}r$ and $w\prec_{N}r^{\prime}$. Therefore, two directed paths $P_{rw}$ and $P_{r^{\prime}w}$ from $r$ respectively $r^{\prime}$ to $w$ exist in $N$. Let $v$ denote the $\preceq_{N}$-maximal vertex contained in $V(P_{rw})\cap V(P_{r^{\prime}w})$. Note that $w\preceq_{N}v\prec_{N}r,r^{\prime}$. Moreover, let $P_{r^{\prime}v}$ denote the directed subpath of $P_{r^{\prime}w}$ from $r^{\prime}$ to $v$. By construction, the four directed paths $P_{rw}$, $P_{r^{\prime}v}$, $P_{w\ell}$ and $P_{w\ell^{\prime}}$ are edge-disjoint. It is now a trivial exercise to show that the subgraph of $N$ consisting of the vertices and edges of these four paths is a subdivision of $\mathsf{X}$ whenever $w=v$ and a subdivision of $\mathsf{X^{\prime}}$ whenever $w\prec_{N}v$. In other words, $N$ contains a $\mathsf{X}$- or $\mathsf{X^{\prime}}$-subdivision $H^{\prime}$ which satisfy $L(H^{\prime})=\{\ell,\ell^{\prime}\}=L(H)$ and $R(H^{\prime})=\{r,r^{\prime}\}=R(H)$.
Assume now that $N$ does not have the global lca-property. By Theorem 3.5, $N$ does not have the global pairwise-lca-property either, i.e., there are distinct vertices $u,v\in V$ such that $|\operatorname{LCA}_{N}(u,v)|\neq 1$. Since $N$ is a network, $\operatorname{LCA}_{N}(\{u,v\})$ is non-empty and we may therefore apply Lemma 4.2 to find a strict $K_{2,2}$-subdivision $H\subseteq N$ that satisfies $R(H)\subseteq\operatorname{LCA}_{N}(\{u,v\})\cap\operatorname{LCA}_{N}(L(H))$. Assume, for contradiction, that there exists an $\mathsf{X}$- or $\mathsf{X^{\prime}}$-subdivision $H^{\prime}\subseteq N$ such that $L(H^{\prime})=L(H)$ and $R(H^{\prime})=R(H)$. One easily verifies that, in $H^{\prime}$, there exists a vertex $w$ such that $\operatorname{LCA}_{H^{\prime}}(L(H^{\prime}))=\{{w}\}$. However, since $R(H^{\prime})=\{r,r^{\prime}\}=R(H)$ and $w\in V(H^{\prime})$ it also holds that ${w}\preceq_{N}r,r^{\prime}$. Since $r$ and $r^{\prime}$ are $\preceq_{N}$-incomparable by assumption, even ${w}\prec_{N}r,r^{\prime}$ holds (cf. Lemma 2.11). In particular, w{w} is a common ancestor of $L(H)=L(H^{\prime})$. Moreover, since $r,r^{\prime}\in\operatorname{LCA}_{N}(L(H))$ the latter arguments show that $w$ is a common ancestor of $L(H)$ strictly below vertices contained in $\operatorname{LCA}_{N}(L(H))$; a contradiction. ∎
5 A constructive characterization
In this section, we provide a way to construct networks with global lca-property starting from the graph $K_{1}$ that consists of a single vertex and no edges. To this end, we show first that every network can be constructed by stepwise adding leaves. We will later refine this process to obtain global lca-networks. To establish this result, we note that a directed graph $G$ is a DAG if and only if it admits a topological order $\lessdot_{G}$ defined on $V(G)$, i.e., the vertices of $G$ can be totally ordered such that $(x,y)\in E(G)$ implies that $x\lessdot_{G}y$ [26].
Lemma 5.1.
Let $\mathbf{N}$ denote the class of all networks. Then, $N\in\mathbf{N}$ if and only if $N\simeq K_{1}$ or $N$ can be constructed from a network $N^{\prime}\in\mathbf{N}$ by applying the following operation:
  1. (O)
    Add a leaf $x\notin V(N^{\prime})$ and edges $(w,x)$ for all $w\in W$, for some non-empty subset $W\subseteq V(N)$.
Proof.
Let $N\in\mathbf{N}$. If $N\simeq K_{1}$, then we are done. Hence, assume that $N$ contains at least two vertices. It is easy to verify that $L(N)\neq\emptyset$ and $\rho_{N}\notin L(N)$ (cf. e.g. [30, p. 6]). Let $x\in L(N)$ and put $W=\operatorname{parent}_{N}(x)$. Consider now the subgraph $N^{\prime}$ of $N$ obtained from $N$ by removal of $x$ and its incident edges. As $x\neq\rho_{N}$ is a leaf and the only vertex that has been removed, $N^{\prime}$ remains a DAG with unique root $\rho_{N}$, i.e., $N^{\prime}\in\mathbf{N}$. Applying Rule (O) to $N^{\prime}$ w.r.t. the set $W$ yields now $N$, which completes the proof of the only-if direction.
For the if direction assume that $N\simeq K_{1}$ or $N$ can be constructed from a network $N^{\prime}\in\mathbf{N}$ by applying Rule (O). Again if $N\simeq K_{1}$, then $N\in\mathbf{N}$. Assume that $N\not\simeq K_{1}$. Since $N^{\prime}$ is, in particular, a DAG, there is a topological order $\lessdot_{N^{\prime}}$ defined on $V(N^{\prime})$. We can now extend this $\lessdot_{N^{\prime}}$ to a topological order $\lessdot_{N}$ defined on $V(N)$ by simply adding $x$ as the $\lessdot_{N}$-maximal element. Thus, $N$ is a DAG [26]. Moreover, $N$ still has only one single root, since the construction of $N$ ensures that $\operatorname{indeg}_{N}(v)=0$ if and only if $v$ coincides with the unique root of $N^{\prime}$. The latter two arguments imply that $N\in\mathbf{N}$. ∎
Suppose that $N$ is obtained from $N^{\prime}$ by applying rule (O), i.e., we add a new leaf $x$ to $N^{\prime}$ and connect $x$ to the vertices in $\emptyset\neq W\subseteq V(N^{\prime})$, then $N^{\prime}=N-x\coloneqq(V(N)\setminus\{x\},E(N)\setminus\{(w,x)\mid w\in W\})$. Hence, it is easy to verify that the following property holds
  1. (P1)
    $u\preceq_{N^{\prime}}v$ if and only if $u\preceq_{N}v$, for all $u,v\in V(N^{\prime})\subseteq V(N)=V(N^{\prime})\cup\{x\}$.
In particular, (P1) implies that the following property is satisfied:
  1. (P2)
    $\operatorname{LCA}_{N}(A)=\operatorname{LCA}_{N^{\prime}}(A)$ for all subsets $A\subseteq V(N^{\prime})$.
Hence, if $N^{\prime}=N-x$ does not have the global lca-property, then there is some subset $A\subseteq V(N^{\prime})$ such that $|\operatorname{LCA}_{N^{\prime}}(A)|\neq 1$. This together with (P2) implies the following
Observation 5.2.
If $x$ is a leaf in a network $N$ and $N^{\prime}=N-x$ does not have the global lca-property, then $N$ does not have the global lca-property.
By Lemma 5.1, every network that has the global lca-property can be constructed starting from a $K_{1}$ and repeated applications of Rule (O). However, not all networks are global lca-networks. Hence, to provide a construction of global lca-networks, we must refine Rule (O) and, in particular, restrict the type of sets $W$ used to add edges $(w,x)$ to $N^{\prime}$. To obtain this result, the following set
$\mathcal{L}_{N^{\prime}}(W|v)\coloneqq\bigcup_{w\in W}\operatorname{LCA}_{{N^{\prime}}}(\{w,v\})$ (2)
defined for networks $N^{\prime}$, a subset $W\subseteq V(N^{\prime})$ and a given vertex $v\in V(N^{\prime})$, will play a central role. In essence, $\mathcal{L}_{N^{\prime}}(W|v)$ is the set comprising all LCAs of $v$ and $w$ for all $w\in W$ in $N^{\prime}$. For example, considering the network $N$ in Figure 3, we have $\mathcal{L}_{N}(\{w_{1},w_{2}\}|w_{3})=\operatorname{LCA}_{N}(\{w_{1},w_{3}\})\cup\operatorname{LCA}_{N}(\{w_{2},w_{3}\})=\{w_{1},u\}$. Note that if $N^{\prime}$ has global lca-property, then $|\operatorname{LCA}_{{N^{\prime}}}(\{w,v\})|=1$ for all $v,w\in V(N^{\prime})$ and Equation (2) simplifies to
$\mathcal{L}_{N^{\prime}}(W|v)\coloneqq\{\operatorname{lca}_{N^{\prime}}(v,w)\,:\,w\in W\}.$ (3)
Lemma 5.3.
Let $N^{\prime}$ be a graph with the global lca-property, and assume that $N$ is obtained from $N^{\prime}$ by applying (O), where $W=\operatorname{parent}_{N}(x)$ denotes the set of parents for the newly added leaf $x\notin V(N^{\prime})$. Then $N$ has the global lca-property if and only if, for all $v\in V(N^{\prime})$, the set $\mathcal{L}_{N^{\prime}}(W|v)$ contains a unique $\preceq_{N^{\prime}}$-minimal vertex.
Proof.
Let $N^{\prime}$ be a network with the global lca-property and suppose that $N$ is obtained from $N^{\prime}$ by application of (O) for some non-empty $W\subseteq V(N^{\prime})$, attaching the new leaf $x$ with the edges $(w,x)$, for each $w\in W$. By construction, we have $W=\operatorname{parent}_{N}(x)$. Note that, by (P2), $\operatorname{LCA}_{N}(A)=\operatorname{LCA}_{N^{\prime}}(A)$ for all subsets $A\subseteq V(N^{\prime})$. Since $N^{\prime}$ has the global lca-property, it follows that, in particular,
  1. (P2enumi)
    $\operatorname{lca}_{N}(u,v)=\operatorname{lca}_{N^{\prime}}(u,v)$ is well-defined for all $u,v\in V(N^{\prime})$.
Before showing the equivalence of the statement of this lemma, we consider the set $\mathcal{L}^{\min}_{N^{\prime}}(W|v)$ that comprises all $\preceq_{N^{\prime}}$-minimal vertices in $\mathcal{L}_{N^{\prime}}(W|v)$ and show that, for every vertex $v\in V(N^{\prime})$, it holds that
$\operatorname{LCA}_{N}(x,v)=\mathcal{L}^{\min}_{N^{\prime}}(W|v),$ (4)
that is, the set of LCAs of $v$ and $x$ in $N$ coincide with the set $\mathcal{L}^{\min}_{N^{\prime}}(W|v)$. Note that $\mathcal{L}_{N^{\prime}}(W|v)$ can be chosen as specified in Equation (3).
Let $v\in V(N^{\prime})=V(N)\setminus\{x\}$ be chosen arbitrarily. We show first that $\operatorname{LCA}_{N}(x,v)\subseteq\mathcal{L}^{\min}_{N^{\prime}}(W|v)$. Note, by Lemma 5.1, $N$ is a network and thus, $\operatorname{LCA}_{N}(x,v)\neq\emptyset$. Hence, we can arbitrarily choose a vertex $u\in\operatorname{LCA}_{N}(x,v)$. Since $x$ is a leaf of $N$ and $v\neq x$, we have $x\prec_{N}u$ and there exists some $p\in\operatorname{parent}_{N}(x)=W$ such that $x\prec_{N}p\preceq_{N}u$. In particular, $u\in V(N^{\prime})$ and (P1) ensures that $v\preceq_{N^{\prime}}u$ and $p\preceq_{N^{\prime}}u$ i.e. $u$ is a common ancestor of $v$ and $p$ in $N^{\prime}$. Since, $N^{\prime}$ has the global lca-property, $\operatorname{lca}_{N^{\prime}}(v,p)$ is well-defined and, by the latter arguments, satisfies $\operatorname{lca}_{N^{\prime}}(v,p)\preceq_{N^{\prime}}u$. Since $u,v,p\in V(N^{\prime})$, we can apply (P1) and (P2enumi) to conclude that $z\coloneqq\operatorname{lca}_{N^{\prime}}(v,p)=\operatorname{lca}_{N}(v,p)\preceq_{N}u$. Since $x\prec_{N}p\preceq_{N}z$ and $v\preceq_{N}z$, the vertex $z$ is a common ancestor of $v$ and $x$ in $N$. Together with $u\in\operatorname{LCA}_{N}(x,v)$ the latter implies that $z\not\prec_{N}u$. Consequently, $u=z=\operatorname{lca}_{N^{\prime}}(v,p)$ and, therefore, $u\in\mathcal{L}_{N^{\prime}}(W|v)$. It remains to show that $u\in\mathcal{L}^{\min}_{N^{\prime}}(W|v)$. Assume, for contradiction, that $u$ is not a $\preceq_{N^{\prime}}$-minimal vertex in $\mathcal{L}_{N^{\prime}}(W|v)$. Hence, there exists some vertex $w\in W$ such that $\operatorname{lca}_{N^{\prime}}(v,w)\prec_{N^{\prime}}u$. Since $u,v,w\in V(N^{\prime})$, we can use (P1) and (P2enumi) to conclude that $\operatorname{lca}_{N^{\prime}}(v,w)=\operatorname{lca}_{N}(v,w)\prec_{N}u$. Since $w\in W=\operatorname{parent}_{N}(x)$, $\operatorname{lca}_{N}(v,w)$ is also a common ancestor of $x$ and $v$ in $N$. This and $\operatorname{lca}_{N}(v,w)\prec_{N}u$ contradicts the fact that $u\in\operatorname{LCA}_{N}(x,v)$. Thus, $u\in\mathcal{L}^{\min}_{N^{\prime}}(W|v)$. As $v\in V(N^{\prime})$ and $u\in{\operatorname{LCA}_{N}(x,v)}$ have been chosen arbitrarily it follows that $\operatorname{LCA}_{N}(x,v)\subseteq\mathcal{L}^{\min}_{N^{\prime}}(W|v)$ for all $v\in V(N^{\prime})$.
We proceed with showing that $\mathcal{L}^{\min}_{N^{\prime}}(W|v)\subseteq\operatorname{LCA}_{N}(x,v)$. Since $N^{\prime}$ has the global lca-property and since $W\neq\emptyset$, it follows that $\mathcal{L}^{\min}_{N^{\prime}}(W|v)\neq\emptyset$. Let $u\in\mathcal{L}^{\min}_{N^{\prime}}(W|v)$ be chosen arbitrarily and let $w\in W$ be a vertex such that $\operatorname{lca}_{N^{\prime}}(v,w)=u$. Since $v,w\in V(N^{\prime})$, (P2enumi) implies that $u=\operatorname{lca}_{N}(v,w)$. Since $w\in W=\operatorname{parent}_{N}(x)$, $u$ is a common ancestor of $x$ and $v$ in $N$. Hence, there is a vertex $u^{\prime}\in\operatorname{LCA}_{N}(x,v)$ such that $u^{\prime}\preceq_{N}u$. To recall, $x$ is a leaf of $N$ and $v\neq x$ which implies that $x\prec_{N}u^{\prime}$ must hold and there exists some parent $p\in W$ of $x$ in $N$ such that $x\prec_{N}p\preceq_{N}u^{\prime}$. This and $u^{\prime}\in\operatorname{LCA}_{N}(x,v)$ implies that $u^{\prime}$ is a common ancestor of $v$ and $p$ in $N$. By (P1), $u^{\prime}$ is a common ancestor of $v$ and $p$ in $N^{\prime}$. Since $N^{\prime}$ has the global lca-property, $\operatorname{lca}_{N^{\prime}}(v,p)$ is well-defined. Taking the latter two arguments together we obtain $\operatorname{lca}_{N^{\prime}}(v,p)\preceq_{N^{\prime}}u^{\prime}\preceq_{N^{\prime}}u$. By construction, $\operatorname{lca}_{N^{\prime}}(v,p)\in\mathcal{L}_{N^{\prime}}(W|v)$. Moreover, since $u\in\mathcal{L}^{\min}_{N^{\prime}}(W|v)$, the vertex $u$ is a $\preceq_{N^{\prime}}$-minimal vertex in $\mathcal{L}_{N^{\prime}}(W|v)$ and thus, $\operatorname{lca}_{N^{\prime}}(v,p)\prec_{N^{\prime}}u$ cannot hold. Taking the latter arguments together we have $\operatorname{lca}_{N^{\prime}}(v,p)=u^{\prime}=u$ which, in turn, together with $u^{\prime}\in\operatorname{LCA}_{N}(x,v)$ implies that $u\in\operatorname{LCA}_{N}(x,v)$. As $u\in\mathcal{L}^{\min}_{N^{\prime}}(W|v)$ was chosen arbitrarily, it follows that $\mathcal{L}^{\min}_{N^{\prime}}(W|v)\subseteq\operatorname{LCA}_{N}(x,v)$. In summary, we have shown that $\mathcal{L}^{\min}_{N^{\prime}}(W|v)=\operatorname{LCA}_{N}(x,v)$ for all $v\in V(N^{\prime})$.
We now return to the main statement. Suppose first that $N$ has the global lca-property. In this case, Theorem 3.5 implies that $N$ has the global pairwise-lca-property. This is, in particular, only possible if $|\operatorname{LCA}_{N}(x,v)|=1$ for all $v\in V(N^{\prime})\subseteq V(N)$. As $\mathcal{L}^{\min}_{N^{\prime}}(W|v)=\operatorname{LCA}_{N}(x,v)$ for all $v\in V(N^{\prime})$ it follows that $|\mathcal{L}^{\min}_{N^{\prime}}(W|v)|=1$ and, therefore, that $\mathcal{L}_{N^{\prime}}(W|v)$ contains a unique $\preceq_{N^{\prime}}$-minimal element for all $v\in V(N^{\prime})$. Conversely, suppose that $\mathcal{L}_{N^{\prime}}(W|v)$ contains a unique $\preceq_{N^{\prime}}$-minimal element for all $v\in V(N^{\prime})$ and thus that $|\mathcal{L}^{\min}_{N^{\prime}}(W|v)|=1$ for all $v\in V(N^{\prime})$. In this case, $\mathcal{L}^{\min}_{N^{\prime}}(W|v)=\operatorname{LCA}_{N}(x,v)$ implies that $|\operatorname{LCA}_{N}(x,v)|=1$ for all $v\in V(N^{\prime})$. Clearly, $|\operatorname{LCA}_{N}(x,x)|=1$. In addition, $|\operatorname{LCA}_{N}(a,b)|=1$ for all $a,b\in V(N^{\prime})$ as $N^{\prime}$ has the global lca-property (c.f. (P2enumi)). The latter three arguments imply that $|\operatorname{LCA}_{N}(a,b)|=1$ for all $a,b\in V(N)$. Thus, $N$ has the global pairwise-lca-property and Theorem 3.5, implies that $N$ has the global lca-property, which completes this proof. ∎
In the following, we show that the class of networks with the global lca-property coincides with the class of holju graphs [^2] We call these graphs holju because the main concept was conceived by some of us while being stranded at Ljubljana Airport (LJU) with a rather significant hangover (HO). We thank the organizers of the 40th TBI-winterseminar in Bled, Slovenia for providing an endless amount of free beer, schnaps and wine, and assure the reader that all proofs in this manuscript were written after a full dose of sleep.
Definition 5.4.
The class of holju graphs is defined recursively. The graph $K_{1}$ is a holju graph. If $G$ is a holju graph, then any graph obtained by applying the following operation once, is also a holju graph:
  1. (Oenumi)
    Add a leaf $x\notin V(G)$ and the edges $(w,x)$ for all $w\in W$, where $W\subseteq V(G)$ is a non-empty subset of vertices such that for any $v\in V(G)$ the set $\mathcal{L}_{G}(W|v)$ contains a unique $\preceq_{G}$-minimal vertex. Note that the case $|W|=1$ is allowed.
Note that (Oenumi) is just a special case of (O). For example, consider the three networks $N$, $N^{\prime}$ and $N^{\prime\prime}$ in Figure 3, where $N=N^{\prime}-x^{\prime}=N^{\prime\prime}-x^{\prime\prime}$. Note that $N$ has the global lca-property, and that both $N^{\prime}$ and $N^{\prime\prime}$ is obtained from $N$ by applying (O), although to different sets $W$. For $N^{\prime}$, we have $W=\{w_{1},w_{2}\}$ and, in particular, $\mathcal{L}_{N}(\{w_{1},w_{2}\}|w_{3})=\{w_{1},u\}$. Since $w_{1}$ and $u$ are $\preceq_{N}$-incomparable, $\mathcal{L}_{N}(\{w_{1},w_{2}\}|w_{3})$ does not contain a unique $\preceq_{N}$-minimal vertex and, consequently, $N^{\prime}$ is not obtained from $N$ by applying (Oenumi). In contrast, $N^{\prime\prime}$ is obtained from $N$ by applying (Oenumi) to $W=\{w_{2},w_{3}\}$, realized by (tediously but with no great effort) verifying that $\mathcal{L}_{N}(W|x)$ indeed has a unique $\preceq_{N}$-minimal vertex for every $x\in V(N)$. In this example, $N^{\prime}$ does not have the global lca-property (since e.g. $|\operatorname{LCA}_{N^{\prime}}(\{x^{\prime},w_{3}\})|=|\{w_{1},u\}|\neq 1$) while $N^{\prime\prime}$ is a global lca-network. This is not a coincidence, due to the following theorem.
Figure 3: Three networks $N$, $N^{\prime}$ and $N^{\prime\prime}$, where $N^{\prime}$ is obtained from $N$ by applying (O) but not (Oenumi), and $N^{\prime\prime}$ is obtained from $N$ by applying (Oenumi). Both NN and $N^{\prime\prime}$ have the global lca-property, while $N^{\prime}$ does not. See the text for more details.
Theorem 5.5.
A network has the global lca-property if and only if it is a holju graph.
Proof.
Let $N$ be a holju graph. Since (Oenumi) is just a special case of (O), Lemma 5.1 ensures $N$ is a network. Note that the constraints imposed on $\mathcal{L}_{G}(W|v)$ in Lemma 5.3 and (Oenumi) are the same. As $N$ is constructed by a sequence of operations of (Oenumi) starting with the $K_{1}$ that trivially has the global lca-property, repeated application of Lemma 5.3, therefore, implies that $N$ has the global lca-property.
Suppose now that $N$ is a network with the global lca-property. If $N\simeq K_{1}$, then $N$ is clearly a holju graph. Suppose that $N\not\simeq K_{1}$. Since $N$ is a network, it can, due to Lemma 5.1, be constructed by stepwise application of (O) starting with a $K_{1}$. This yields a sequence of constructed networks $K_{1}=N_{1},\,N_{2},\,\dots,\,N_{k}=N$, where $k>1$. Let us denote with $x_{i}$ the new leaf added to $N_{i-1}$ and with $W_{i}$ the set of parents of $x_{i}$ in $N_{i}$ used in the construction of $N_{i}$ when applying (O) to $N_{i-1}$, for each $i\in\{2,\dots,k\}$. In other words, $N_{i-1}=N_{i}-x_{i}$ for the leaf $x_{i}$ of $N_{i}$, for each $i\in\{2,\dots,k\}$. Since $N=N_{k}$ has the global lca-property, Observation 5.2 implies that $N_{k-1}$ must have the global lca-property. By induction, $N_{i}$ must have the global lca-property, $1\leq i\leq k$. Since each $N_{i}$ is obtained from $N_{i-1}$ by applying (O) and since $N_{i-1}$ has the global lca-property, Lemma 5.3 implies that $\mathcal{L}_{N_{i-1}}(W_{i}|v)$ contains a unique $\preceq_{N_{i-1}}$-minimal vertex, for each $i\in\{2,\dots,k\}$. Hence, to obtain $N_{i}$ from $N_{i-1}$ we have, in fact, applied (Oenumi), $i\in\{2,\dots,k\}$. Therefore, each $N_{i}$ and thus, $N$ is a holju graph. ∎
6 A characterization by tree-leaf-child and leaf-extended DAGs
A tree-leaf $v$ in a DAG $G$ is a leaf $v\in L(G)$ with in-degree one. A DAG is tree-leaf-child if every inner vertex is adjacent to a tree-leaf. Tree-leaf-child DAGs form a subclass of the tree-child DAGs, in which every inner vertex is adjacent to vertex with with in-degree one, but not necessarily to a leaf [8]. For later reference we provide
Lemma 6.1 ([22, Cor. 20]).
A tree-child network $G$ is an lca-network if and only if its clustering system $\mathfrak{C}_{N}$ is closed.
Lemma 6.2.
Every tree-leaf-child DAG $G$ satisfies (PCC).
Proof.
To see that $G$ satisfies (PCC), let $u,v\in V(G)$ be two distinct vertices. If $u\preceq_{G}v$, then $\operatorname{\mathtt{C}}_{G}(u)\subseteq\operatorname{\mathtt{C}}_{G}(v)$ by Lemma 2.4. Assume now that $\operatorname{\mathtt{C}}_{G}(u)\subseteq\operatorname{\mathtt{C}}_{G}(v)$. If $u$ is a leaf in $G$, then $\operatorname{\mathtt{C}}_{G}(u)=\{u\}$ and thus, $u\in\operatorname{\mathtt{C}}_{G}(v)$. By definition, $u\preceq_{G}v$. Assume now that $u$ is an inner vertex in $G$. Since $G$ is tree-leaf-child, there is a leaf $x_{u}$ whose unique parent is $u$. Moreover, $x_{u}\in\operatorname{\mathtt{C}}_{G}(v)$. Thus, there is a directed path $P$ from $v$ to $x_{u}$. Taking the latter arguments together, this path $P$ must contain $u$. Hence, $u\preceq_{G}v$. In summary, $G$ satisfies (PCC). ∎
Figure 4: Two networks NN and $N^{\prime}$, neither of which have the global (pairwise)-lca-property. Moreover, NN does have the pairwise-lca-property i.e. $\operatorname{lca}_{N}(A)$ is well-defined for each subset $A\subseteq L(N)$ of size $|A|=2$, yet $\operatorname{LCA}_{N}(\{a,b,c\})=\{w,w^{\prime}\}$ and $N$ therefore does not have the lca-property. In $N^{\prime}$, there is a unique LCA for every non-empty subset of leaves, but $\operatorname{LCA}_{N^{\prime}}(\{u,v\})=\{w,w^{\prime}\}$, implying that $N^{\prime}$ is not a global lca-network. Moreover, both $N$ and $N^{\prime}$ are networks in which every inner vertex is adjacent to a leaf, indicating the need to enforce leaf-tree-child in Proposition 6.4. Lastly, for $N$, the set system $\mathfrak{C}_{N}$ is a pre-binary clustering system that is not closed since $\operatorname{\mathtt{C}}_{N}(w)\cap\operatorname{\mathtt{C}}_{N}(w^{\prime})=\{a,b,c\}\notin\mathfrak{C}_{N}$. The latter argument shows that the Statements (5) and (6) in Proposition 6.4 are, in the general setting, not equivalent.
In general DAGs, the pairwise-lca-property and the lca-property do not imply knowledge about the global lca-property nor of the global pairwise-lca-property. To see this consider the network $N$ and $N^{\prime}$ in Figure 4. Here, $\operatorname{lca}_{N}(A)$ is well-defined for each subset $A\subseteq L(N)$ of size $|A|=2$ and $N$ consequently has the pairwise-lca-property. Despite this, it does not have the lca-property (and therefore not the global lca-property) since, for example, $|\operatorname{LCA}_{N}(\{a,b,c\})|=|\{w,w^{\prime}\}|\neq 1$. In the network $N^{\prime}$ of the same figure, $\operatorname{lca}_{N^{\prime}}(A)$ is well-defined for each non-empty $A\subseteq L(N^{\prime})$ and therefore $N^{\prime}$ has both the lca-property and the pairwise-lca-property, yet $|\operatorname{LCA}_{N^{\prime}}(u,v)|=|\{w,w^{\prime}\}|\neq 1$ and $N^{\prime}$ does not have the global lca-property. Nevertheless, we can connect the global (pairwise-)lca-property with the (pairwise-)lca-property in the case of tree-leaf-child DAGs. To this end we will need to consider the following property of clustering systems:
Definition 6.3.
[5] A clustering system $\mathfrak{C}$ is pre-binary if for every $x,y\in X$ there is unique inclusion-minimal element $C\in\mathfrak{C}$ that contains $x$ and $y$.
We are now in the position to extend Lemma 6.1 in the following manner:
Proposition 6.4.
The following statements are equivalent for every tree-leaf-child DAG $G$:
  1. (1)
    $G$ has the global lca-property.
  2. (2)
    $G$ has the global pairwise-lca-property.
  3. (3)
    $G$ has the lca-property.
  4. (4)
    $G$ has the pairwise-lca-property.
  5. (5)
    $\mathfrak{C}_{G}$ is a closed clustering system.
  6. (6)
    $\mathfrak{C}_{G}$ is a pre-binary clustering system.
Proof.
Let $G$ be a tree-leaf-child DAG on $X$. By Theorem 3.5, Condition (1) and (2) are equivalent. Moreover, Condition (1) clearly implies (3), and Condition (3) implies (4). We continue with showing that Condition (4) implies (2) to establish the equivalence between Conditions (1), (2), (3) and (4).
To this end, we assume, by contraposition, that $G$ does not have the global pairwise-lca-property. In this case, there are distinct vertices $x,y\in V(G)$ such that $|\operatorname{LCA}_{G}(\{x,y\})|\neq 1$. If $x,y\in L(G)$, then it immediately follows that $G$ does not have the pairwise-lca-property. Hence consider the case when at least one of $x$ and $y$ is an inner vertex of $G$. Without loss of generality, assume that $x\notin X$. Since $G$ is tree-leaf-child, we can find a tree-leaf $x^{\prime}$ such that $x^{\prime}$ is a child of $x$. Similarly, if $y\notin X$ we can find a tree-leaf $y^{\prime}$ such that $y^{\prime}$ is a child of $y$. If $y$ is a leaf itself, put $y^{\prime}=y$. Since $x^{\prime}\preceq_{G}x$ and $y^{\prime}\preceq_{G}y$, every common ancestor of $x$ and $y$ is also a common ancestor of $x^{\prime}$ and $y^{\prime}$. Moreover, since $x^{\prime}$ is a tree-leaf and a child of $x$, every ancestors of $x^{\prime}$ – except $x^{\prime}$ itself – is also an ancestor of $x$. If $y^{\prime}\prec_{G}y$, analogous arguments shows that every ancestors of $y^{\prime}$ – except $y^{\prime}$ itself – is also an ancestor of $y$ and, if $y=y^{\prime}$, the ancestors of $y$ and $y^{\prime}$ clearly coincide. In summary so-far, we have
$\mathtt{ANC}_{G}(\{x,y\})\subseteq\mathtt{ANC}_{G}(\{x^{\prime},y^{\prime}\})\subseteq\mathtt{ANC}_{G}(\{x,y\})\cup\{x^{\prime},y^{\prime}\}.$
Note that $|\operatorname{LCA}_{G}(\{x,y\})|\neq 1$ implies that neither $x\preceq_{G}y$ nor $y\preceq_{G}x$ can hold. Together with $x\neq y$, this implies that $x$ and $y$ are $\preceq_{G}$-incomparable. Since $x^{\prime}$ has as unique parent $x$ and either $y=y^{\prime}$ or $y^{\prime}$ has as unique parent $y$, it follows that $x^{\prime}$ and $y^{\prime}$ must be $\preceq_{G}$-incomparable. It is now easy to see that each vertex that is a common ancestor of $x^{\prime}$ and $y^{\prime}$ is a common ancestor of $x$ and $y$, i.e., $\mathtt{ANC}_{G}(\{x,y\})=\mathtt{ANC}_{G}(\{x^{\prime},y^{\prime}\})$. In particular $\operatorname{LCA}_{G}(\{x,y\})=\operatorname{LCA}_{G}(\{x^{\prime},y^{\prime}\})$ and, consequently, $G$ does not have the pairwise-lca-property.
Thus, Conditions (1), (2), (3) and (4) are equivalent. Moreover, since $G$ is a tree-leaf-child DAG, Lemma 6.2 implies that $G$ satisfied (PCC). This allows us to use Proposition 3.4, to conclude that (1) and (5) are equivalent.
To complete the proof, we show the equivalence between Condition (4) and (6). First observe that Theorem 3.6 of [21] states that if $H$ is a DAG that satisfy (PCC), then $H$ has the pairwise-lca-property if and only if $\mathfrak{C}_{H}$ is pre-binary. As Lemma 6.2 ensures that $G$ satisfy (PCC), the latter argument immediately implies that $G$ satisfies Condition (4) if and only if it satisfies (6). ∎
The class of tree-leaf-child DAGs is of course quite restrictive. Nevertheless, we can associate a type of canonical tree-leaf-child DAG $\operatorname{lxt}(G)$ to every DAG $G$ which will be used later to characterize DAGs with global lca-property.
Definition 6.5.
For every DAG $G=(V,E)$, the corresponding leaf-extended DAG $\operatorname{lxt}(G)=(V^{*},E^{*})$ is obtained by putting first $V^{*}=V$ and $E^{*}=E$ and then, adding, for every inner vertex $v\in V$, a new vertex $x_{v}$ to $V^{*}$ and the edge $(v,x_{v})$ to $E^{*}$.
We collect now a couple of results for leaf-extended DAGs. Note that, by definition, $\operatorname{lxt}(G)$ is tree-leaf-child for every DAG $G$.
Lemma 6.6.
Let $G$ be a DAG and $\operatorname{lxt}(G)$ be its leaf-extended version. Then, $\operatorname{outdeg}_{\operatorname{lxt}(G)}(v)>1$ for all inner vertices in $\operatorname{lxt}(G)$. Moreover, $\operatorname{lxt}(G)$ satisfies (PCC) and $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(u)\neq\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)$ for all distinct $u,v\in V(\operatorname{lxt}(G))$. In particular, $\operatorname{lxt}(G)$ is lca-relevant and $\operatorname{sf}(\operatorname{lxt}(G))$ is regular.
Proof.
Let $G$ be a DAG and $G^{*}\coloneqq\operatorname{lxt}(G)$ be its leaf-extended version. By construction, for each inner vertex $v$ in $G^{*}$ there is a unique leaf $x_{v}$ whose unique parent is $v$ and such that $x_{v}\notin V(G)$. In particular, $v$ is an inner vertex in $G$. Hence, $\operatorname{outdeg}_{G}(v)\geq 1$ and thus, $\operatorname{outdeg}_{G^{*}}(v)=\operatorname{outdeg}_{G}(v)+1>1$. Since $G^{*}$ is tree-leaf-child, Lemma 6.2 implies that $G^{*}$ satisfies (PCC).
To show that $G^{*}$ is lca-relevant, we equivalently show that the condition in Theorem 2.12(4) is satisfied. Suppose that $\operatorname{\mathtt{C}}_{G^{*}}(u)\subseteq\operatorname{\mathtt{C}}_{G^{*}}(v)$ for some vertices $u,v\in V(G^{*})$. If $u$ is a leaf of $G^{*}$, then $\operatorname{\mathtt{C}}_{G^{*}}(u)=\{u\}$ and $u\in\operatorname{\mathtt{C}}_{G^{*}}(v)$ immediately implies that $u\preceq_{G^{*}}v$. If $u$ is an inner vertex of $G^{*}$ then, since $G^{*}$ is tree-leaf-child, $u$ has a leaf-child $x_{u}$ such that $u$ is the unique parent of $x_{u}$ in $G^{*}$. Clearly, $x_{u}\in\operatorname{\mathtt{C}}_{G^{*}}(u)$ and, consequently, $x_{u}\in\operatorname{\mathtt{C}}_{G^{*}}(v)$. In particular any directed path from $v$ to $x_{u}$ must contain $u$, and such a path must exist by $x_{u}\in\operatorname{\mathtt{C}}_{G^{*}}(v)$, from which $u\preceq_{G^{*}}v$ follows. The latter together with Lemma 2.4 implies that Theorem 2.12(4) holds and that $G^{*}$ is lca-relevant. Moreover, Theorem 2.12(5) implies that $\operatorname{sf}(G^{*})$ is regular. Finally, Theorem 2.12(3) ensures that $\operatorname{\mathtt{C}}_{G^{*}}(u)\neq\operatorname{\mathtt{C}}_{G^{*}}(v)$ for all distinct $u,v\in V(G^{*})$. ∎
Note that tree-leaf-child DAGs $G$ may contain inner vertices with out-degree one and distinct vertices $u$ and $v$ with $\operatorname{\mathtt{C}}_{G}(u)=\operatorname{\mathtt{C}}_{G}(v)$ (e.g. a rooted edge). Hence Lemma 6.6 does, in general, not hold for arbitrary tree-leaf-child DAGs.
Theorem 6.7.
The following statements are equivalent for every DAG $G$ on $X$:
  1. (1)
    $G$ has the global lca-property.
  2. (2)
    $\operatorname{lxt}(G)$ has the global lca-property.
  3. (3)
    $\operatorname{lxt}(G)$ has the global pairwise-lca-property.
  4. (4)
    $\operatorname{lxt}(G)$ has the lca-property.
  5. (5)
    $\operatorname{lxt}(G)$ has the pairwise-lca-property.
  6. (6)
    $\mathfrak{C}_{\operatorname{lxt}(G)}$ is a closed clustering system.
  7. (7)
    $\mathfrak{C}_{\operatorname{lxt}(G)}$ is a pre-binary clustering system.
Proof.
Since $\operatorname{lxt}(G)$ is tree-leaf-child, Proposition 6.4 implies that Conditions (2)–(7) are equivalent. Therefore, we only need to establish the equivalence between Condition (1) and (2). Suppose that the DAG $G$ has the global lca-property. By definition, $\operatorname{lxt}(G)$ coincides with the DAG obtained from $G$ by using operation (O) for all subsets $W=\{w\}$ with $w\in V^{0}(G)$. Since $G$ has the global lca-property, $|\operatorname{LCA}_{G}(\{v,w\})|=1$ for all $v\in V(G)$ and $w\in V^{0}(G)$. Hence, for all $w\in V^{0}(G)$, the set $\mathcal{L}_{G}(\{w\}|v)=\bigcup_{w^{\prime}\in\{w\}}\operatorname{LCA}_{G}(\{w^{\prime},v\})=\operatorname{LCA}_{G}(\{w,v\})$ as defined in Equation (2) has a unique element and, thus, in particular, a unique $\preceq_{G}$-minimal vertex. Theorem 5.5 implies that $G$ is a holju graph. The latter arguments together with Definition 5.4 imply that $\operatorname{lxt}(G)$ is a holju graph. Again, Theorem 5.5 implies that $\operatorname{lxt}(G)$ has the global lca-property. Hence, Condition (1) implies (2).
For the converse, assume that $G$ does not have the global lca-property, and let $A\subseteq V(G)$ be a non-empty subset of vertices such that $|\operatorname{LCA}_{G}(A)|\neq 1$. By construction, $A\subseteq V(G)\subseteq V(\operatorname{lxt}(G))$ and $\mathtt{ANC}_{\operatorname{lxt}(G)}(A)=\mathtt{ANC}_{G}(A)$. Therefore, $|\operatorname{LCA}_{\operatorname{lxt}(G)}(A)|\neq 1$ must hold, and $\operatorname{lxt}(G)$ does not have the global lca-property. By contraposition, Condition (2) implies (1). ∎
Prominent examples of phylogenetic networks in the literature are galled-trees [19] and, more general, level-1 networks [38]. A network $N$ is a galled-tree if each maximal biconnected subgraph $K$ in $N$ is either a single vertex, an edge or there are two vertices $u$ and $v$ in $K$ such that $K$ is composed of exactly two directed $uv$-paths that only have $u$ and $v$ in common [22]. In particular, galled-trees form a subclass of level-1 networks, i.e., networks $N$ in which each biconnected component contains at most one vertex $v$ with $\operatorname{indeg}_{N}(v)>1$.
Corollary 6.8.
Every level-1 network and thus every galled-tree is a global lca-network and thus a holju graph.
Proof.
Let $N=(V,E)$ be a level-1 network. It is easy to see that $\operatorname{lxt}(N)$ remains a level-1 network. As shown in Lemma 49 in [22] every level-1 network is an lca-network. Thus, $\operatorname{lxt}(N)$ is an lca-network. By Theorem 6.7, $N$ is a global lca-network. By Theorem 5.5, $N$ is a holju graph. Since galled-trees are level-1 networks, in particular galled-trees are global lca-network and holju graphs. ∎
The characterization in Proposition 6.4 and, via leaf-extension, in Theorem 6.7, shows that for tree-leaf-child DAGs the global lca-property is equivalent to the corresponding cluster system being pre-binary, and hence equivalent to closedness. It is therefore natural to place these results into the broader context of pre-$k$-ary set systems and closure theory. The next two paragraphs briefly recall this connection and explain why the closedness conditions appearing above can be viewed as a special case of a more general set-system framework.
In [9] a property denoted by (kKC) was considered that was defined as follows: A set system $\mathfrak{C}$ is (kKC) if, for every non-empty $A\subseteq V$ with $|A|\leq k$ there is a unique inclusion-minimal $C\in\mathfrak{C}$ that contains $A$. Note that, for $k=2$, this reduces to $\mathfrak{C}$ being pre-binary. This condition can be expressed equivalently in terms of the closure function
$\operatorname{cl}(A)\coloneqq\bigcap_{\begin{subarray}{c}C\in\mathfrak{C}\\ A\subseteq C\end{subarray}}C$ (5)
as follows: For all non-empty $A\subseteq V$ with $|A|\leq k$ it holds that $A\subseteq\operatorname{cl}(A)\in\mathfrak{C}$. That is, $\operatorname{cl}(A)$ is the unique inclusion-minimal $C\in\mathfrak{C}$ that contains $A$ appearing in the definition of (kKC). Set systems with property (kKC) were called pre-$k$-ary in [39]. As a generalization, [21] considers pre-$\mathscr{I}$-ary set systems, where the condition $|A|\leq k$ is replaced by $|A|\in\mathscr{I}$ for some set $\mathscr{I}$ of positive integers. Hence, $\mathfrak{C}$ is pre-$|V|$-ary in the language of [39] if and only if it is pre-$\mathscr{I}$-ary for $\mathscr{I}=\{1,\dots,|V|\}$. Theorem 3.6 in [21] shows that a DAG $G$ that satisfies (PCC) has the lca-property if and only if $\mathfrak{C}_{G}$ it is pre-$|V|$-ary.
For pre-$|V|$-ary set systems, $\operatorname{cl}(A)$ is well-defined in Equation (5) for all non-empty sets $A\subseteq V$ and coincides with the conventional definition of the closure function. In particular, in this case, $A\in\mathfrak{C}$ if and only if $A=\operatorname{cl}(A)$ and $A\neq\emptyset$. There is a subtle – albeit trivial – difference between the definition of the closure function in the literature on convex set systems [40] and in the setting of clustering systems: Since the empty set does not have an interpretation as a cluster, we define $\emptyset\notin\mathfrak{C}$, while in convexities one usually stipulates $\emptyset\in\mathfrak{C}$. Clustering systems therefore are considered closed if they are closed under non-empty intersections, and we have $A\in\mathfrak{C}$ if and only $A=\operatorname{cl}(A)$ and $A\neq\emptyset$. In the setting of convexities, we have $A\cap B\in\mathfrak{C}$ for all $A,B\in\mathfrak{C}$ and equivalently $A\in\mathfrak{C}$ iff $A=\operatorname{cl}(A)$. The definition of pre-$|V|$-ary thus coincides with $\mathfrak{C}$ being a closed set system, i.e., an abstract convexity, up to the exclusion of the empty set. Clearly there is a 1-1 correspondence between set systems that include and exclude the empty set, respectively, and definitions such as the ones of closure and closedness above are easily adjusted to the respective setting. A brief discussion of this point can also be found in [9].
7 Characterization via descendants, ancestors, and intermediaries
Recall that the DAG $\operatorname{lxt}(G)$ is obtained from a DAG $G$ by adding, for all inner vertices $v$ in $G$ a new tree-leaf $x_{v}$ whose unique parent in $\operatorname{lxt}(G)$ is $v$. Hence, $x_{v}$ is such a new leaf in $\operatorname{lxt}(G)$ if and only if $x_{v}\in\overline{L}(\operatorname{lxt}(G))\coloneqq L(\operatorname{lxt}(G))\setminus L(G)$. The following technical results is used to establish further characterizations of global lca-DAGs. In particular, we show how the descendant system $\mathfrak{D}_{G}$ can be obtained from $\mathfrak{C}_{\operatorname{lxt}(G)}$ and $L(G)$ and that $(\mathfrak{C}^{*},\subseteq)$, $(\mathfrak{D}_{G},\subseteq)$, and $(V(G),\preceq_{G})$ are pairwise isomorphic, where $\mathfrak{C}^{*}$ is obtained from $\mathfrak{D}_{G}$ by removal of all singletons.
Lemma 7.1.
For a DAG $G$, the set system $\mathfrak{D}_{G}$ can be obtained from $\mathfrak{C}_{\operatorname{lxt}(G)}$ by applying the following steps
  1. (1)
    remove the singleton $\{x_{v}\}$ from $\mathfrak{C}_{\operatorname{lxt}(G)}$ for all $x_{v}\in\overline{L}\coloneqq L(\operatorname{lxt}(G))\setminus L(G)$ and, afterwards,
  2. (2)
    replace all $x_{v}\in\overline{L}\cap C$ in all $C\in\mathfrak{C}_{\operatorname{lxt}(G)}$ by the respective vertex $v\in V(G)$.
Moreover, if $\mathfrak{C}^{*}$ is the set obtained from $\mathfrak{C}_{\operatorname{lxt}(G)}$ after application of Step (1), then there is a 1-to-1 correspondence between the elements in $\mathfrak{C}^{*}$ and the elements in $\mathfrak{D}_{G}$. Furthermore, the posets $(\mathfrak{C}^{*},\subseteq)$, $(\mathfrak{D}_{G},\subseteq)$ and $(V(G),\preceq_{G})$ are isomorphic.
Proof.
Let $G$ be a DAG and $\overline{L}\coloneqq L(\operatorname{lxt}(G))\setminus L(G)$. Since $L(G)\subseteq L(\operatorname{lxt}(G))$, we have $\mathtt{D}_{G}(x)=\{x\}\in\mathfrak{D}_{G}$ for all $x\in L(G)$. Furthermore, $x_{v}\notin L(G)$ for all $x_{v}\in\overline{L}$ and it follows that $\{x_{v}\}\notin\mathfrak{D}_{G}$ for all $x_{v}\in\overline{L}$. Consequently, the set of all singleton sets in $\mathfrak{D}_{G}$ can be obtained from the set of all singleton sets in $\mathfrak{C}_{\operatorname{lxt}(G)}$ by applying Step (1).
Assume now that Step (1) has been applied to $\mathfrak{C}_{\operatorname{lxt}(G)}$ and call the resulting set $\mathfrak{C}^{*}$. We proceed with showing that $D\coloneqq\mathtt{D}_{G}(v)\in\mathfrak{D}_{G}$ with $|D|>1$ if and only if $D$ can be obtained from $C\coloneqq\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)\in\mathfrak{C}^{*}$ by applying Step (2). In other words, all non-singleton sets in $\mathfrak{D}_{G}$ can be obtained from all non-singleton sets in $\mathfrak{C}^{*}$ by applying Step (2).
Suppose that $w\in C$ and thus, that $w\in L(\operatorname{lxt}(G))$ and $w\preceq_{\operatorname{lxt}(G)}v$. Assume that $w\in\overline{L}$, then $w=x_{z}$ for some $z\in V(G)$. In this case, $w$ will be replaced by $z$ in $C$. This step is correct, as in this case, $x_{z}\preceq_{\operatorname{lxt}(G)}v$ implies that $z\preceq_{G}v$ and thus, $z\in D$. Assume now that $w\notin\overline{L}$ and thus, that $w\in L(G)$. Then, $w\preceq_{\operatorname{lxt}(G)}v$ and thus, $w\preceq_{G}v$. As Step (2) is not applied for $w$ in this case, $w$ remains in $D$. Hence, the vertices in $C$ to which Step (2) has been applied yield a subset of vertices in $D$.
Suppose that $w\in D$ and thus that $w\preceq_{G}v$. Assume first that $w\in L(G)\subseteq L(\operatorname{lxt}(G))$. Hence, $w\preceq_{G}v$ and $w\preceq_{\operatorname{lxt}(G)}v$. This and the fact that $w\notin\overline{L}$ implies that Step (2) is not applied to $w$, i.e., $w\in C$. Assume now that $w\in V^{0}(G)$. In this case, $w\notin L(\operatorname{lxt}(G))$, which implies that $w\notin C$. However, since $w$ is an inner vertex, there is a tree-leaf-child $x_{w}\in\overline{L}$ with $x_{w}\prec_{\operatorname{lxt}(G)}w$. Moreover, $w\preceq_{G}v$ implies $w\preceq_{\operatorname{lxt}(G)}v$ and thus, $x_{w}\in C$. In the latter case, Step (2) is applied and $x_{w}$ is replaced by $w$.
In summary, each vertex in $D$ is either already contained in $C$ or the result of application of Step (2). Therefore, $D\coloneqq\mathtt{D}_{G}(v)\in\mathfrak{D}_{G}$ with $|D|>1$ if and only if $D$ can be obtained from $C\coloneqq\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)\in\mathfrak{C}^{*}$ by applying Step (2). Hence, $\mathfrak{D}_{G}$ can be obtained from $\mathfrak{C}_{\operatorname{lxt}(G)}$ by first applying Step (1) and then Step (2).
As shown above, each singleton $\{x\}\in\mathfrak{C}^{*}$ corresponds to the unique singleton $\{x\}\in\mathfrak{D}_{G}$ (in which case $x\in L(G)$) and each $\mathtt{D}_{G}(v)\in\mathfrak{D}_{G}$ with $|\mathtt{D}_{G}(v)|>1$ is obtained from $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)\in\mathfrak{C}^{*}$ by application of Step (2). By Lemma 2.6, $\mathtt{D}_{G}(u)\neq\mathtt{D}_{G}(v)$ for all distinct $u,v\in V(G)$ and, by Lemma 6.6, $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(u)\neq\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)$ for all distinct $u,v\in V(\operatorname{lxt}(G))$. Summarizing the latter arguments, there is a 1-to-1 correspondence between the elements in $\mathfrak{C}^{*}$ and the elements in $\mathfrak{D}_{G}$.
We continue with showing that $(\mathfrak{D}_{G},\subseteq)$ and $(\mathfrak{C}^{*},\subseteq)$ are isomorphic. Based on the latter arguments, the map $\varphi\colon\mathfrak{D}_{G}\to\mathfrak{C}^{*}$ where $\mathtt{D}_{G}(v)\in\mathfrak{D}_{G}$ is mapped to $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)\in\mathfrak{C}^{*}$ for all $v\in V(G)$ is bijective. Let $\mathtt{D}_{G}(v),\mathtt{D}_{G}(w)\in\mathfrak{D}_{G}$. By Lemma 2.6, $\mathtt{D}_{G}(v)=\mathtt{D}_{G}(w)$ if and only if $v=w$. By Lemma 6.6, $v=w$ if and only if $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)=\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(w)$. Suppose now that $\mathtt{D}_{G}(v)\subsetneq\mathtt{D}_{G}(w)$ and thus, that $v\neq w$. Lemma 2.6 implies that $v\prec_{G}w$. By construction, $v,w\in V(\operatorname{lxt}(G))$ and $v\prec_{\operatorname{lxt}(G)}w$. This together with Lemma 2.4 and 6.6 implies that $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)\subsetneq\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(w)$. Conversely, suppose that $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)\subsetneq\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(w)$ and thus that $v\neq w$. By Lemma 6.6, $\operatorname{lxt}(G)$ satisfies (PCC). However, $w\preceq_{G}v$ cannot holds because in this case Lemma 2.4 would imply that $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(w)\subseteq\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(v)$. Hence, $v\prec_{G}w$ must hold. Lemma 2.6 implies that $\mathtt{D}_{G}(v)\subsetneq\mathtt{D}_{G}(w)$. In summary, $(\mathfrak{C}^{*},\subseteq)$ and $(\mathfrak{D}_{G},\subseteq)$ are isomorphic. Since, by Lemma 2.6, $(\mathfrak{D}_{G},\subseteq)$ and $(V(G),\preceq_{G})$ are isomorphic, we conclude that $(\mathfrak{C}^{*},\subseteq)$ and $(V(G),\preceq_{G})$ are isomorphic. ∎
A similar result can be obtained to show that $\mathfrak{C}_{\operatorname{lxt}(G)}$ can be reconstructed from $\mathfrak{D}_{G}$ by applying the inverse of Step (2), and then the inverse of Step (1).
Lemma 7.2.
Let $G$ be a DAG. If $\mathfrak{C}_{\operatorname{lxt}(G)}$ is closed, then $\mathfrak{D}_{G}$ is closed.
Proof.
Let $G$ be a DAG and assume that $\mathfrak{C}_{\operatorname{lxt}(G)}$ is closed. We denote with $\mathfrak{C}^{*}$ the set that is obtained from $\mathfrak{C}_{\operatorname{lxt}(G)}$ after application of Step (1) in Lemma 7.1, i.e., after removal of all singletons $\{x_{v}\}$ from $\mathfrak{C}_{\operatorname{lxt}(G)}$, where $x_{v}\in\overline{L}\coloneqq L(\operatorname{lxt}(G))\setminus L(G)$. Let $A,B\in\mathfrak{D}_{G}$ be such that $A\cap B\neq\emptyset$. By Lemma 7.1, there is a 1-to-1 correspondence between the elements in $\mathfrak{D}_{G}$ and $\mathfrak{C}^{*}$. In particular, there is a unique element $A^{\prime}\in\mathfrak{C}^{*}$ from which $A$ has been obtained after application of Step (2) on $A^{\prime}$. In other words, all $x_{v}$ in $\overline{L}\cap A^{\prime}$ have been replaced in $A^{\prime}$ by the respective vertex $v\in V(G)$ to obtain $A$. Similarly, there is a unique $B^{\prime}\in\mathfrak{C}^{*}$ from which $B$ has been derived after application of Step (2) on $B^{\prime}$. In particular, any $x_{v}\in A^{\prime}\cap B^{\prime}$ with $x_{v}\in\overline{L}$ has been replaced by $v$ in $A$ and $B$ while all other vertices $w\in A^{\prime}\cap B^{\prime}$ with $w\notin\overline{L}$ remain in $A$ and $B$ and therefore $C^{\prime}\coloneqq A^{\prime}\cap B^{\prime}\neq\emptyset$.
Since $A^{\prime},B^{\prime}\in\mathfrak{C}^{*}\subseteq\mathfrak{C}_{\operatorname{lxt}(G)}$ and $\mathfrak{C}_{\operatorname{lxt}(G)}$ is closed, it follows that $C^{\prime}\in\mathfrak{C}_{\operatorname{lxt}(G)}$. Assume, for contradiction, $C^{\prime}\notin\mathfrak{C}^{*}$. In this case, $C^{\prime}=\{x_{v}\}$ for some $x_{v}\in\overline{L}$. Since $A^{\prime},B^{\prime}\in\mathfrak{C}^{*}$, it must hold that $A^{\prime}\neq\{x_{v}\}$ and $B^{\prime}\neq\{x_{v}\}$. Consequently, the vertex $v$ to which $x_{v}$ was attached is an inner vertex in $G$, and there is a leaf $z\neq x_{v}$ with $z\in L(G)\subseteq L(\operatorname{lxt}(G))$ such that $z\prec_{\operatorname{lxt}(G)}v$. Moreover, there are vertices $u,w$ in $\operatorname{lxt}(G)$ for which $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(u)=A^{\prime}$ and $\operatorname{\mathtt{C}}_{\operatorname{lxt}(G)}(w)=B^{\prime}$. Since $x_{v}\in A^{\prime},B^{\prime}$ and since $x_{v}$ has as its unique parent $v$, it follows that $v\preceq_{\operatorname{lxt}(G)}u,w$ and thus, that $z\prec_{\operatorname{lxt}(G)}u,w$. Therefore, $z\in A^{\prime}$ and $z\in B^{\prime}$ and, in particular, $z\in C^{\prime}$. Since $z\neq x_{v}$ it follows that $C^{\prime}\neq\{x_{v}\}$; a contradiction. Consequently, $C^{\prime}\in\mathfrak{C}^{*}$ must hold. Due to the 1-to-1 correspondence between the elements in $\mathfrak{D}_{G}$ and $\mathfrak{C}^{*}$ we have an element $C\in\mathfrak{D}_{G}$ that is obtained from $C^{\prime}$ by replacing the vertices in $C^{\prime}$ according to Step (2) in Lemma 7.1. This in particular implies that $C=A\cap B$. Since $C\in\mathfrak{D}_{G}$, the set $\mathfrak{D}_{G}$ is closed. ∎
Let us now turn to sets $\mathtt{ANC}_{G}(A)$ of all common ancestors of $A$ in $G$. In the following we denote with $\mathfrak{A}_{G}$ the set comprising $\mathtt{ANC}_{G}(\{v\})$ for all $v\in V(G)$ in a DAG $G$. By slight abuse of notation we write $\mathtt{ANC}_{G}(v)$ instead of $\mathtt{ANC}_{G}(\{v\})$ when $\mathtt{ANC}_{G}$ is applied to singleton sets.
Observation 7.3.
Let $G$ be a DAG. Then, for all vertices $v\in V$ it holds that $\mathtt{D}_{G}(v)\cap\mathtt{ANC}_{G}(v)=\{v\}$. Moreover, every descendant (resp., ancestor) of $u$ in $G$ is an ancestor (resp. descendant) of $u$ in $\operatorname{rev}(G)$ and, therefore, $\mathfrak{A}_{G}=\mathfrak{D}_{\operatorname{rev}(G)}$ and $\mathfrak{D}_{G}=\mathfrak{A}_{\operatorname{rev}(G)}$.
Lemma 7.4.
Let $G=(V,E)$ be a DAG such that $\mathfrak{D}_{G}$ is closed. Then, for all $x,y\in V(G)$ with $\mathtt{ANC}_{G}(\{x,y\})\neq\emptyset$ it holds that $|\operatorname{LCA}_{G}(\{x,y\})|=1$.
Proof.
Let $G=(V,E)$ be a DAG such that $\mathfrak{D}_{G}$ is closed and let $x,y\in V$ be such that $\mathtt{ANC}\coloneqq\mathtt{ANC}_{G}(\{x,y\})\neq\emptyset$. Note that any common ancestor of $x$ and $y$ must be contained in $\mathtt{ANC}$. Since $\mathtt{ANC}\neq\emptyset$, we have $\operatorname{LCA}_{G}(\{x,y\})\neq\emptyset$. Assume, for contradiction, that $|\operatorname{LCA}_{G}(\{x,y\})|>1$ and let $w,w^{\prime}\in\operatorname{LCA}_{G}(\{x,y\})$ be distinct. By definition of LCAs, the elements in $\operatorname{LCA}_{G}(\{x,y\})$ must be pairwise $\preceq_{G}$-incomparable. By definition, $x,y\in D\coloneqq\mathtt{D}_{G}(w)\cap\mathtt{D}_{G}(w^{\prime})$. Since $\mathfrak{D}_{G}$ is closed, $D\in\mathfrak{D}_{G}$ and, hence, there is a vertex $v\in V$ such that $\mathtt{D}_{G}(v)=D$. Since, by definition, $v\in\mathtt{D}_{G}(v)$, it follows that $v$ is a descendant of $w$ and $w^{\prime}$ and thus, $v\preceq_{G}w,w^{\prime}$. This together with the fact that all elements in $\operatorname{LCA}_{G}(\{x,y\})$ are pairwise $\preceq_{G}$-incomparable implies that $v\notin\operatorname{LCA}_{G}(\{x,y\})$ and thus, $v\prec_{G}w,w^{\prime}$; a contradiction to $w,w^{\prime}\in\operatorname{LCA}_{G}(\{x,y\})$. Hence, $|\operatorname{LCA}_{G}(\{x,y\})|=1$ must hold. ∎
Lemma 7.5.
Let $G$ be a DAG. Then, $\mathfrak{D}_{G}$ is closed if and only if $\mathfrak{A}_{G}$ is closed.
Proof.
Let $G$ be a DAG. Assume first that $\mathfrak{D}_{G}$ is closed. Let $A,B\in\mathfrak{A}_{G}$ be such that $A\cap B\neq\emptyset$. Note that, by definition, there are vertices $u,v\in V(G)$ with with $A=\mathtt{ANC}_{G}(u)$ and $B=\mathtt{ANC}_{G}(v)$. In particular, $\mathtt{ANC}_{G}(\{u,v\})=A\cap B\neq\emptyset$. Hence, we can apply Lemma 7.4 to conclude that $\operatorname{LCA}_{G}(\{u,v\})=\{w\}$ for some $w\in V(G)$. By definition, $w\preceq_{G}z$ for all $z\in\mathtt{ANC}_{G}(\{u,v\})$ which implies that $\mathtt{ANC}_{G}(\{u,v\})\subseteq\mathtt{ANC}_{G}(w)$. Moreover, $u,v\preceq_{G}w$ implies that every $z\in\mathtt{ANC}_{G}(w)$ satisfies $u,v\preceq_{G}z$ and, therefore, $z\in\mathtt{ANC}_{G}(\{u,v\})$. Hence, $\mathtt{ANC}_{G}(w)\subseteq\mathtt{ANC}_{G}(\{u,v\})$. In summary, $\mathtt{ANC}_{G}(\{u,v\})=A\cap B=\mathtt{ANC}_{G}(w)\in\mathfrak{A}_{G}$. Thus, $\mathfrak{A}_{G}$ is closed.
For the converse, assume that $\mathfrak{A}_{G}$ is closed. By Observation 7.3, $\mathfrak{A}_{G}=\mathfrak{D}_{\operatorname{rev}(G)}$. Hence, $\mathfrak{D}_{\operatorname{rev}(G)}$ is closed. Since $\operatorname{rev}(G)$ is a DAG, we can re-use the previous arguments to conclude that $\mathfrak{A}_{\operatorname{rev}(G)}$ is closed. Since, by Observation 7.3, $\mathfrak{A}_{\operatorname{rev}(G)}=\mathfrak{D}_{G}$ it follows that $\mathfrak{D}_{G}$ is closed. ∎
Definition 7.6.
For a DAG $G$ and two of its vertices $u,v\in V(G)$ let $\operatorname{\mathtt{B}}_{G}(u,v)\coloneqq\mathtt{D}_{G}(u)\cap\mathtt{ANC}_{G}(v)$ be set of all intermediary vertices located between $u$ and $v$, i.e., the set of all vertices $w\in V(G)$ satisfying $v\preceq_{G}w\preceq_{G}u$. The intermediary system is then $\mathfrak{I}_{G}\coloneqq\{\operatorname{\mathtt{B}}_{G}(u,v)\mid u,v\in V(G)\}$.
It is easy to see that $\operatorname{\mathtt{B}}_{G}(u,v)$ contains all vertices that lie on any directed $uv$-path. The map $(u,v)\mapsto\operatorname{\mathtt{B}}_{G}(u,v)$ coincides with the directed all-path transit function of the DAG, see [4] for more details. Note that, in a DAG $G$, it must hold that $\operatorname{\mathtt{B}}_{G}(v,v)=\{v\}$ for all $v\in V(G)$. Hence, $\{v\}\in\mathfrak{I}_{G}$ for all $v\in V(G)$. However, $\mathfrak{I}_{G}$ is in general not grounded since, for any two vertices $u$ and $v$ for which $u\prec_{G}v$ holds or where $u$ is $\preceq_{G}$-incomparable to $v$, we have $\operatorname{\mathtt{B}}_{G}(u,v)=\emptyset$ and therefore, $\emptyset\in\mathfrak{I}_{G}$. Hence, $\mathfrak{I}_{G}$ is grounded if and only if it contains a single vertex only. Moreover, $\mathfrak{I}_{G}$ is not necessarily rooted.
Lemma 7.7.
Let $N$ be a network, then $\mathfrak{A}_{N}$ is closed if and only if $\mathfrak{I}_{N}$ is closed.
Proof.
Let $N$ be a network. Assume first that $\mathfrak{A}_{N}$ is closed. By Lemma 7.5, $\mathfrak{D}_{N}$ is closed. Suppose there are vertices $x,y,u,v\in V(N)$ such that $\mathtt{B}_{N}(x,y)\cap\mathtt{B}_{N}(u,v)\neq\emptyset$. We obtain
$\emptyset\neq\mathtt{B}_{N}(x,y)\cap\mathtt{B}_{N}(u,v)=\ \mathtt{D}_{N}(x)\cap\mathtt{ANC}_{N}(y)\cap\mathtt{D}_{N}(u)\cap\mathtt{ANC}_{N}(v)=\$$\underbrace{\mathtt{D}_{N}(x)\cap\mathtt{D}_{N}(u)}_{\neq\emptyset}\cap\underbrace{\mathtt{ANC}_{N}(y)\cap\mathtt{ANC}_{N}(v)}_{\neq\emptyset}$
[…]
where we have used that $\mathfrak{A}_{N}$ and $\mathfrak{D}_{N}$ are closed to justify the existence of vertices $w,z\in V(G)$ that satisfy $\mathtt{D}_{N}(w)=\mathtt{D}_{N}(x)\cap\mathtt{D}_{N}(u)$ and $\mathtt{ANC}_{N}(z)=\mathtt{ANC}_{N}(y)\cap\mathtt{ANC}_{N}(v)$. Hence, if $\mathfrak{A}_{N}$ is closed then $\mathfrak{I}_{N}$ is closed.
Now suppose that $\mathfrak{I}_{N}$ is closed. Note that for every $v\in V(N)$, we have $\mathtt{ANC}_{N}(v)=\mathtt{B}_{N}(\rho,v)$, where $\rho$ denotes the unique root of $N$. Hence, for every $x,y\in V(N)$ such that $\mathtt{ANC}_{N}(x)\cap\mathtt{ANC}_{N}(y)\neq\emptyset$, it holds that $\mathtt{B}_{N}(\rho,x)\cap\mathtt{B}_{N}(\rho,y)\neq\emptyset$. Thus closedness of $\mathfrak{I}_{N}$ implies the existence of $w,z\in V(N)$ such that $\mathtt{B}_{N}(w,z)=\mathtt{B}_{N}(\rho,x)\cap\mathtt{B}_{N}(\rho,y)$. In particular, $\rho\in\mathtt{B}_{N}(w,z)$ holds. Since $\mathtt{B}_{N}(w,z)=\mathtt{D}_{N}(w)\cap\mathtt{ANC}_{N}(z)$, we have $\rho\in\mathtt{D}_{N}(w)$. This and the fact that $\rho$ is the unique $\preceq_{N}$-maximal vertex in $N$ implies that $w=\rho$, that is, $\mathtt{B}_{N}(w,z)=\mathtt{B}_{N}(\rho,z)=\mathtt{ANC}_{N}(z)$. Summarizing, we have
$\mathtt{ANC}_{N}(x)\cap\mathtt{ANC}_{N}(y)=\mathtt{B}_{N}(\rho,x)\cap\mathtt{B}_{N}(\rho,y)=\mathtt{B}_{N}(\rho,z)=\mathtt{ANC}_{N}(z)\in\mathfrak{A}_{N},$
and $\mathfrak{A}_{N}$ is closed. ∎
Theorem 7.8.
The following statements are equivalent for every network $N$.
  1. (1)
    NN has the global lca-property.
  2. (2)
    $\mathfrak{D}_{N}$ is closed.
  3. (3)
    $\mathfrak{A}_{N}$ is closed.
  4. (4)
    $\mathfrak{I}_{N}$ is closed.
Proof.
If $N$ has the global lca-property, then Theorem 6.7 implies that $\mathfrak{C}_{\operatorname{lxt}(N)}$ is a closed clustering system. By Lemma 7.2, $\mathfrak{D}_{N}$ is closed. Suppose that $\mathfrak{D}_{N}$ is closed. Since $N$ is a network, $\mathtt{ANC}_{N}(\{x,y\})\neq\emptyset$ for all $x,y\in V(N)$. This and Lemma 7.4 implies that $|\operatorname{LCA}_{G}(x,y)|=1$ for all $x,y\in V(N)$. Hence, $N$ has the global pairwise-lca-property and thus, by Theorem 3.5, the global lca-property. In summary (1) and (2) are equivalent. By Lemma 7.5, (2) and (3) are equivalent and by Lemma 7.7, (3) and (4) are equivalent. ∎
8 The Canonical Hasse-diagram of a DAG
A common theme throughout this contribution is that the existence of unique least common ancestors is closely tied to order-theoretic structures. For example, networks $N$ with global lca-property are precisely those DAGs whose underlying reachability poset $(V(N),\preceq_{N})$ forms a join-semilattice (Theorem 3.9). Moreover, several of our characterizations are expressed in terms of Hasse diagrams of set systems associated with a DAG. From this perspective, it is natural to ask to what extent a DAG can be recovered from such Hasse-diagram representations.
Intuitively, the Hasse diagrams of set systems derived from a DAG $G$ should reflect the partial order $(V(G),\preceq_{G})$ that governs ancestor relations and, hence, the structure of least common ancestors. This is particularly true for the poset $(\mathfrak{D}_{G},\subseteq)$, which is isomorphic to $(V(G),\preceq_{G})$ by Lemma 2.6. As an immediate consequence of Lemmas 2.6 and 2.8, we therefore have $\operatorname{sf}(G)\simeq\mathscr{H}(\mathfrak{D}_{G},\subseteq)$. This relationship extends to the Hasse diagram of the poset $(\mathfrak{C}^{*},\subseteq)$ obtained from $\mathfrak{C}_{\operatorname{lxt}(G)}$, due to its isomorphism with $(\mathfrak{D}_{G},\subseteq)$ established in Lemma 7.1. In contrast, the set system $\mathfrak{C}_{G}$, although directly defined in terms of descendant leaves and hence closely related to LCAs, often does not provide sufficient information to reconstruct $G$ (cf., e.g., Figure 1). This situation improves substantially after passing to the leaf-extended DAG $\operatorname{lxt}(G)$: by Theorem 6.7, the LCA-properties of $G$ are faithfully reflected in $\operatorname{lxt}(G)$, while Lemma 7.1 shows that the cluster system $\mathfrak{C}_{\operatorname{lxt}(G)}$ encodes the descendant system $\mathfrak{D}_{G}$ and thus, the poset $(\mathfrak{D}_{G},\subseteq)$, which is isomorphic to $(V(G),\preceq_{G})$. Hence, the poset $(\mathfrak{C}_{\operatorname{lxt}(G)},\subseteq)$ yields a canonical order-theoretic representation of the reachability structure underlying the LCA structure of $G$.
The aim of this section is to make this relationship explicit. More precisely, we show that, although $\mathfrak{C}_{G}$ alone is generally insufficient to recover $G$, the Hasse diagram of the poset $(\mathfrak{C}_{\operatorname{lxt}(G)},\subseteq)$ contains enough information to reconstruct $G$ up to shortcuts after a simple canonical pruning operation. Thus, this section complements the preceding LCA characterizations by showing how the order structure underlying least common ancestors can be recovered from an associated set of clusters.
Definition 8.1.
A $G$-lopped DAG $H$ is obtained form a tree-leaf-child DAG $G$ by removal of exactly one leaf-child $x_{v}$ of $v$ that has in-degree one, for all inner vertices $v\in V(H)$. We denote by $\operatorname{lop}(G)$ an arbitrary $G$-lopped DAG.
Lemma 8.2.
Let $G$ be a tree-leaf-child DAG. If $H$ and $H^{\prime}$ are two $G$-lopped DAGs, then $H\simeq H^{\prime}$. Moreover, $\operatorname{sf}(\operatorname{lop}(G))\simeq\operatorname{lop}(\operatorname{sf}(G))$.
Proof.
Let $G=(V,E)$ be a tree-leaf-child DAG on $X$ and suppose that $H$ and $H^{\prime}$ are two $G$-lopped DAGs. If $H=H^{\prime}$, there is nothing to show. Assume that $H\neq H^{\prime}$. In this case, $H$ is obtained from $G$ by removing a tree-leaf $x_{v}\in X$ and $H^{\prime}$ is obtained from $G$ by removing a tree-leaf $x^{\prime}_{v}\in X$ with $x_{v}\neq x_{v}^{\prime}$ for at least one inner vertex $v$ of $G$. We call such particular inner vertices $v$ $(H,H^{\prime})$-leaf-deviating. Note that all $(H,H^{\prime})$-leaf-deviating vertices $v$ satisfy $v\in V(H)\cap V(H^{\prime})$ as they must be non-leaf vertices. We can now define a bijective map $\varphi\colon V(H)\to V(H^{\prime})$ by putting $\varphi(v)=v$ if $v\in V(H^{\prime})$ and $\varphi(x_{v})=x^{\prime}_{v}$ for all $(H,H^{\prime})$-leaf-deviating vertices $v\in V(H)$. In particular, the restriction of $\varphi$ to $V^{0}(G)\subseteq V(H)\cap V(H^{\prime})$ is just the identity map. It is now an easy exercise to verify that $\varphi$ is an isomorphism between $H$ and $H^{\prime}$.
We continue with proving that $\operatorname{sf}(\operatorname{lop}(G))\simeq\operatorname{lop}(\operatorname{sf}(G))$. Note that if there is a shortcut $e=(u,v)$ in $G$ for which the vertex $v$ is a leaf, then $v$ must be a hybrid in $G$ but may become a tree-leaf in $\operatorname{sf}(G)$ and thus, might be removed when considering $\operatorname{lop}(\operatorname{sf}(G))$. To address this issue, we mark, for each inner vertex $v$ in $G$, exactly one tree-leaf-child of $v$ red and call it $x_{v}$. The latter is possible since $G$ is tree-leaf-child. Let $S\subseteq E$ be the set of all shortcuts in $G$. The DAG $\operatorname{sf}(G)$ is obtained from $G$ by removal of all edges in $S$. Lemma 2.3 implies that $\operatorname{sf}(G)$ is $\preceq_{G}$-preserving, i.e., $u\preceq_{G}v$ precisely if $u\preceq_{\operatorname{sf}(G)}v$ for all $u,v\in V$. Thus, $X=L(\operatorname{sf}(G))$ and $V^{0}\coloneqq V^{0}(G)=V^{0}(\operatorname{sf}(G))$. Moreover, the in-degree of a vertex $w$ in $\operatorname{sf}(G)$ cannot be higher than the respective in-degree of $w$ in $G$, i.e., tree-leafs of $G$ remain tree-leafs in $\operatorname{sf}(G)$. Hence, all red-marked vertices in $G$ specify, in particular, a unique tree-leaf-child $x_{v}$ in $\operatorname{sf}(G)$ for all $v\in V^{0}$. We can thus obtain a particular $\operatorname{sf}(G)$-lopped DAG, denoted $\operatorname{lop}_{\text{\tiny red}}(\operatorname{sf}(G))$, by removing all red-marked vertices from $\operatorname{sf}(G)$. Similarly, we obtain a particular $G$-lopped DAG, denoted $\operatorname{lop}_{\text{\tiny red}}(G)$, by removing all red-marked vertices from $G$. It is easy to verify that the set of shortcuts in $\operatorname{lop}_{\text{\tiny red}}(G)$ coincides with $S$. Hence, to obtain $\operatorname{sf}(\operatorname{lop}_{\text{\tiny red}}(G))$ we can simply remove all edges in $S$. As we have applied identical operations to obtain $\operatorname{sf}(\operatorname{lop}_{\text{\tiny red}}(G))$ as well as $\operatorname{lop}_{\text{\tiny red}}(\operatorname{sf}(G))$ from $G$ it follows that $\operatorname{sf}(\operatorname{lop}_{\text{\tiny red}}(G))=\operatorname{lop}_{\text{\tiny red}}(\operatorname{sf}(G))$. Now, as argued in the previous paragraph, all $G$-lopped DAGs are isomorphic. Hence, $\operatorname{lop}(G)\simeq\operatorname{lop}_{\text{\tiny red}}(G)$ for any $\operatorname{lop}(G)$. In addition, $\operatorname{sf}(H)$ is uniquely determined for any DAG $H$ which together with $\operatorname{lop}(G)\simeq\operatorname{lop}_{\text{\tiny red}}(G)$ implies that $\operatorname{sf}(\operatorname{lop}(G))\simeq\operatorname{sf}(\operatorname{lop}_{\text{\tiny red}}(G))$. By analogous arguments, it holds that $\operatorname{lop}(\operatorname{sf}(G))\simeq\operatorname{lop}_{\text{\tiny red}}(\operatorname{sf}(G))$ for every possible $\operatorname{lop}(\operatorname{sf}(G))$. Taking the latter two arguments together with $\operatorname{sf}(\operatorname{lop}_{\text{\tiny red}}(G))=\operatorname{lop}_{\text{\tiny red}}(\operatorname{sf}(G))$ we can conclude that $\operatorname{sf}(\operatorname{lop}(G))\simeq\operatorname{lop}(\operatorname{sf}(G))$ always holds. ∎
We close this section by providing Proposition 8.3, which shows that the order-theoretic structure underlying ancestor relations—and hence the structure from which least common ancestors arise—can be canonically recovered from the cluster system of the leaf-extended DAG $\operatorname{lxt}(G)$, even when the original cluster system $\mathfrak{C}_{G}$ is insufficient for this purpose.
Proposition 8.3.
Let $G$ be a DAG and $\operatorname{lxt}(G)$ its leaf-extended version. Then, $\operatorname{sf}(G)\simeq\operatorname{lop}(\mathscr{H}(\mathfrak{C}_{\operatorname{lxt}(G)},\subseteq))$. In other words, every DAG $G$ is isomorphic to a graph obtained from $\operatorname{lop}(\mathscr{H}(\mathfrak{C}_{\operatorname{lxt}(G)},\subseteq))$ by adding a possibly empty set of shortcuts.
Proof.
Set $\mathscr{H}\coloneqq\mathscr{H}(\mathfrak{C}_{\operatorname{lxt}(G)},\subseteq)$. By Lemma 6.6, $\operatorname{sf}(\operatorname{lxt}(G))$ is regular. Thus, $\operatorname{sf}(\operatorname{lxt}(G))\simeq\mathscr{H}(\mathfrak{C}_{\operatorname{sf}(\operatorname{lxt}(G))},\subseteq)$. Note that $\operatorname{sf}(\operatorname{lxt}(G))$ is uniquely determined and that, by Lemma 2.3, $\mathfrak{C}_{\operatorname{sf}(\operatorname{lxt}(G))}=\mathfrak{C}_{\operatorname{lxt}(G)}$. Consequently, $\operatorname{sf}(\operatorname{lxt}(G))\simeq\mathscr{H}$. By Lemma 8.2, $G\simeq\operatorname{lop}(\operatorname{lxt}(G))$. Hence, $\operatorname{sf}(G)\simeq\operatorname{sf}(\operatorname{lop}(\operatorname{lxt}(G)))\simeq\operatorname{lop}(\operatorname{sf}(\operatorname{lxt}(G)))\simeq\operatorname{lop}(\mathscr{H})$, and $G$ is thus isomorphic to a DAG obtained from $\operatorname{lop}(\mathscr{H})$ by adding some $k\geq 0$ shortcuts. ∎
9 Concluding Remarks
In this contribution, we have characterized global lca-networks or, equivalently, global lca-DAGs, i.e., those DAGs that admit a unique LCA for every subset of their vertices. It does not come as a big surprise that global lca-DAGs $G$ are precisely those DAGs where the underlying poset $(V(G),\preceq_{G})$ is a join-semilattice. However, as shown here, global lca-DAGs (and therefore join-semilattices) exhibit additional rich structural properties: they are characterized by certain topological minors, can be constructed stepwise from a single-vertex graph, and are precisely those DAGs where the LCA is uniquely defined for each pair of vertices. In addition, global lca-DAGs can be recognized in polynomial time and are exactly those networks $N$ for which their descendant-system $\mathfrak{D}_{N}$ (resp., ancestor-system $\mathfrak{A}_{N}$ or intermediary system $\mathfrak{I}_{N}$) is closed. Finally, we have shown that each DAG $G$ can be uniquely reconstructed from the Hasse diagram of the clustering system $\mathfrak{C}_{\operatorname{lxt}(G)}$ of its leaf-extended version $\operatorname{lxt}(G)$.
The characterization of global lca-DAGs in terms of holju graphs, i.e., the stepwise construction of such DAGs from a single-vertex graph, is of particular interest. As shown in Lemma 5.1, every network $N=(V,E)$ can be constructed from a single-vertex graph by stepwise addition of a new vertex $x$, which is connected to some subset $W\subseteq V$. The characterization of global lca-DAGs then involves restricting this set $W$ in each of the construction steps. It is of interest to understand in more detail which DAGs with specified properties admit a similar construction. In particular, how does the addition of $x$, linked to the vertices in $W$, influence properties such as the existence of LCAs or clustering systems with desired properties? For instance, the clustering system of a level-1 network has been characterized as those set systems that are closed and satisfy a property called (L) [22]. In particular, level-1 networks are holju graphs (cf. Cor 6.8). This raises the question of which properties of $W$ can be relaxed to ensure only that $\mathfrak{C}_{G}$ is closed. Similarly, what conditions must be imposed on $W$ so that the resulting network is an lca-network but not necessarily a global lca-network?
In contrast to clustering systems, the closedness of the underlying descendant system $\mathfrak{D}_{N}$, ancestor system $\mathfrak{A}_{N}$, and intermediary system $\mathfrak{I}_{N}$ characterizes the global lca-property. Clearly, these three set systems contain all the information about the underlying network needed to reconstruct it (up to its potential shortcuts) and thus are nearly as informative as the underlying adjacency matrix. In practical applications, one is often faced with the problem that only partial information about the underlying set systems is provided. For example, consider a (shortcut-free) DAG $G$ and its descendant system $\mathfrak{D}_{G}$. Now consider a set system $\mathfrak{F}$ such that each $F\in\mathfrak{F}$ satisfies $F\subseteq D$ for some $D\in\mathfrak{D}_{G}$, denoted by $\mathfrak{F}\subseteq\mathfrak{D}_{G}$. In other words, $\mathfrak{F}$ is obtained from $\mathfrak{D}_{G}$ by removing some elements from certain $D\in\mathfrak{D}_{G}$. In this case, $\mathfrak{F}$ contains only partial information about $\mathfrak{D}_{G}$. This raises the question: to what extent can $G$ be recovered from $\mathfrak{F}$? While this question might be difficult to answer in general, it may be feasible for DAGs equipped with structural constraints (e.g., lca-networks or global lca-networks). The latter question is also is closely related to the question of how much information about $G$ is contained in the “containment-closure” $\mathrm{CCL}(\mathfrak{F})=\cap_{H,\mathfrak{F}\subseteq\mathfrak{D}_{H}}\mathfrak{D}_{H}$, which is the intersection of all $\mathfrak{D}_{H}$ over all DAGs $H$ for which $\mathfrak{F}\subseteq\mathfrak{D}_{H}$ holds. A similar concept has been studied for so-called triples in phylogenetic trees [18].
The semi-lattice convexity is known to be a convex geometry where every convex set has the extreme-point property [40], i.e., each convex set $K$ contains a point $k\in K$ such that $K\setminus\{k\}$ is again convex. Every convex set is therefore the convex hull of its extreme points. In our case, the extreme points of the join-semilattice $(D,\sqsubseteq)$ are precisely the least common ancestors of the join-subsemilattices $(A,\sqsubseteq),\ A\subseteq D$ that are again convex. The co-points of an element $p$ of this convex geometry are maximal convex join-subsemilattices that do not contain the element $p$. It is of practical interest for network reconstruction to study to what extend $(D,\sqsubseteq)$ can be reconstructed from co-points of certain restricted number of leaves.
Funding
All authors declare that they have no conflicts of interest.
This work was supported in part by the DST, Govt. of India (Grant No. DST/INT/DAAD/P-03/ 2023 (G)), the DAAD, Germany (Grant No. 57683501), and the CSIR-HRDG for the Senior Research Fellowship (09/0102(12336)/2021-EMR-I). PFS acknowledges the financial support by the Federal Ministry of Education and Research of Germany (BMBF) through DAAD project 57616814 (SECAI, School of Embedded Composite AI), and jointly with the Sächsische Staatsministerium für Wissenschaft, Kultur und Tourismus in the programme Center of Excellence for AI-research Center for Scalable Data Analytics and Artificial Intelligence Dresden/Leipzig, project identification number: SCADS24B.